### v2

# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

Abstract
We introduce a practical circuit-complexity model for fixed-precision neural networks and apply it to explain (and partially overcome) a persistent learnability barrier in impartial games, with NIM as the canonical example. Our model captures bounded-depth, polynomial-size, fixed-precision neural inference (including finite-window recurrent and attention-style models) and shows that these models can be simulated by $\mathrm{AC}^{0}$ circuits. This places them below $\mathrm{TC}^{0}$ and, in particular, below exact parity and exact majority capabilities.
On the negative side, we prove that single-frame AlphaZero-style agents with such $\mathrm{AC}^{0}$-constrained evaluation/policy networks cannot achieve strong mastery of NIM (optimal play from every reachable position), even when augmented with any polynomial-time bounded search. The core obstruction is the necessity of global parity (nim-sum) information.
On the positive side, we show that augmenting the state with a short window of recent history (two frames suffice for NIM) exposes locally computable transition information. In particular, nimber differences between consecutive positions are $\mathrm{AC}^{0}$-computable when the positions differ in a constant number of heaps. This supports a simple local restoration response rule: after an opponent move, one can restore the zero nim-sum invariant by matching the observed nimber difference, without ever recomputing the global nim-sum from scratch.
Empirically, we validate the restoration viewpoint in a controlled supervised setting: from teacher–random transitions in 20-heap, 4-bit NIM, the two-frame policy exhibits a two-stage learning curve (an intermediate plateau around 0.54 followed by near-perfect restoration accuracy), while the one-frame baseline stays near chance. When combined with bounded per-move MCTS, the learned 2F priors translate into a large head-to-head advantage (Section 6.2).
Finally, we address an important practical question: why $\mathrm{AC}^{0}$ is a good model for feasible learnability even though practitioners observe that neural networks can learn “majority-like” behaviour. The resolution is that both $\mathrm{AC}^{0}$ and realistic fixed-precision networks can learn and compute approximate majority very well, while the sharpness and reliability needed to leverage majority gates to compute parity (as in $\mathrm{TC}^{0}$ constructions) is not feasible under fixed precision and noisy training/inference.
1 Introduction
AlphaZero-style reinforcement learning has achieved striking successes across a range of perfect-information games [18, 17, 15, 16]. Motivated by Harvey Friedman’s proposal of NIM as an AI challenge [8], a series of empirical studies explored whether “standard” AlphaZero pipelines can master impartial games such as NIM [24, 22, 25]. A recurring phenomenon is that, despite the apparent simplicity of NIM, learning optimal play is blocked by global parity constraints (the nim-sum).
This paper develops a theory that both (i) explains why this barrier is robust under standard single-frame training and (ii) shows how a representation change (multi-frame history) enables optimal play without requiring more powerful neural computation.
1.1 Contributions
  • An $\mathrm{AC}^{0}$ model for fixed-precision neural inference. We formalize three “weak” neural models (feedforward threshold networks, finite-window recurrent networks, and finite-window attention-style models) with constant depth, polynomial size, and fixed precision weights/thresholds. We show that they can be simulated by $\mathrm{AC}^{0}$ circuits (Theorem 2.2), with a clean construction in Appendix B.
  • Single-frame impossibility for strong mastery. We prove that $\mathrm{AC}^{0}$-constrained single-frame policy/evaluation networks cannot strongly master NIM, and polynomial-time bounded search cannot overcome this (Section 4.2).
  • Multi-frame possibility via local delta and restoration. With two-frame (or bounded) history, we show that local nimber-difference computations are $\mathrm{AC}^{0}$-computable (Lemma 4.5) and that a player can restore the zero nim-sum invariant after an opponent move by matching the observed nimber difference (Proposition 4.6). This bypasses the need to compute the global nim-sum from scratch on every move.
  • Empirical validation via supervised restoration and search. We instantiate the restoration principle as a supervised learning problem using teacher–random transitions: the input is either the current position (1F) or the last two positions (2F), and the label is a (possibly non-unique) optimal restoration/winning move from the post-opponent position. In a 20-heap, 4-bit NIM setting with a dataset of $10^{6}$ examples, the one-frame model remains at a near-random baseline (about $8\%$ restoration accuracy), while the two-frame model exhibits a two-stage learning curve: it first rises to an intermediate plateau around $0.54$ and later breaks through to essentially perfect restoration accuracy. When paired with per-move MCTS (fixed simulation budget), the learned 2F priors yield a large head-to-head Elo advantage over 1F under the same search budget (Section 6.2).
  • Why $\mathrm{AC}^{0}$ is a good practical learnability model. We clarify the apparent tension that “majority feels feasible” in practice while parity remains infeasible: what is learnable/feasible is approximate majority, which is compatible with $\mathrm{AC}^{0}$, whereas using majority to compute parity requires a level of sharpness/precision that is not feasible under fixed precision and noise (Section 5).
1.2 Paper organisation
Section 2 introduces the fixed-precision neural models and the $\mathrm{AC}^{0}$ simulation theorem. Section 3 recalls impartial games and NIM. Section 4 presents the impossibility and possibility results. Section 5 discusses learnability and the $\mathrm{AC}^{0}$ vs $\mathrm{TC}^{0}$ modelling issue. Section 6 reports experiments, including a supervised restoration benchmark that sharply separates one-frame and two-frame policies (Section 6.2). Appendices provide formal definitions and full circuit constructions.
2 Theoretical framework and neural network models
2.1 Fixed-precision neural models and $\mathrm{AC}^{0}$ simulation
We consider neural inference models that are (i) bounded depth, (ii) polynomial size, and (iii) fixed precision: all weights and thresholds are drawn from a finite set of representable values that does not grow with the input length. Concretely, we fix constants:
  • a precision parameter (e.g. bit-width) captured by a fixed denominator $D\in\mathbb{N}$,
  • a magnitude bound $W\in\mathbb{N}$,
  • and (for recurrent/attention models) a finite time window $T\in\mathbb{N}$.
We define
$\mathcal{Q}_{W,D}\;=\;\Big\{\frac{k}{D}\ :\ k\in\{0,1,2,\dots,W\}\Big\},$
a finite set independent of input length.
Each neuron computes a thresholded weighted sum of Boolean inputs. Formal definitions for $\mathsf{NN}$, $\mathsf{RNN}$ and $\mathsf{LTST}$ are in Appendix A.
Definition 2.1 (Neural networks with constant precision (informal)).
A fixed-precision $\mathsf{NN}$ is a constant-depth feedforward network of polynomially many threshold neurons, whose weights and thresholds lie in $\mathcal{Q}_{W,D}$. Analogously, fixed-precision $\mathsf{RNN}$ and $\mathsf{LTST}$ models have constant depth and polynomial size per time step and operate within a fixed time window $T$.
Why $\mathrm{AC}^{0}$ and not $\mathrm{TC}^{0}$? (Key modelling paragraph)
A common first reaction is to model constant-depth threshold networks by $\mathrm{TC}^{0}$, since $\mathrm{TC}^{0}$ includes majority/threshold gates. However, the definition of $\mathrm{TC}^{0}$ (as used in circuit complexity) implicitly allows thresholds and weights whose description length can scale with input length (typically $\Theta(\log n)$ bits, or more generally polylogarithmic), enabling exact majority and other sharp global counting predicates. In contrast, our focus is fixed precision neural inference: every real parameter is stored in a fixed word size (e.g. 32-bit), so the set of representable thresholds and weights is finite and does not scale with $n$. Under this constraint, the “knife-edge” sharpness required for exact majority on arbitrarily large inputs is not available; what is feasible in practice is learning approximate majority behaviour (high accuracy away from the threshold), which is already achievable within $\mathrm{AC}^{0}$. Crucially, the degree of sharpness needed to use majority gates as reliable building blocks for parity computations is not feasible under fixed precision and noise; this explains why parity remains unlearnable/unstable even though majority-like behaviour is learnable (expanded in Section 5).
2.2 Simulation by circuits
We state a single theorem that applies to all three models.
Theorem 2.2 (Simulation of fixed-precision neural models by $\mathrm{AC}^{0}$).
Fix constants $W,D\in\mathbb{N}$ and a time window $T\in\mathbb{N}$. Let $\{\mathcal{M}_{n}\}_{n\geq 1}$ be any family of models of type $\mathsf{NN}$, $\mathsf{RNN}$ (with window $T$), or $\mathsf{LTST}$ (with window $T$) such that:
  1. (1)
    (Polynomial size) the number of neurons/units in $\mathcal{M}_{n}$ is at most $\mathrm{poly}(n)$.
  2. (2)
    (Constant depth) the number of layers is a fixed constant $L$ (independent of $n$); for $\mathsf{RNN}/\mathsf{LTST}$ the unrolled computation uses at most $T$ time steps.
  3. (3)
    (Fixed precision) all weights and thresholds lie in $\mathcal{Q}_{W,D}$.
Then the Boolean function computed by $\mathcal{M}_{n}$ can be computed by an $\mathrm{AC}^{0}$ circuit family of polynomial size and constant depth (depending only on $L,T,W,D$).
3 Impartial games and NIM
We briefly recall standard definitions; see [2, 3, 4, 5] for extensive background.
3.1 Impartial games
Definition 3.1 (Impartial game).
An impartial game is a pair $(\mathcal{P},\mathcal{M})$ where $\mathcal{P}$ is a nonempty set of positions and $\mathcal{M}(p)\subseteq\mathcal{P}$ is the (finite) set of legal moves from $p$, with the property that the move relation does not depend on which player is to play (symmetry). We assume normal play: the player who makes the last move wins.
A position with no legal moves is terminal.
3.2 NIM
Definition 3.2 (NIM).
A NIM position consists of $h$ heaps with sizes $x_{1},\dots,x_{h}\in\mathbb{N}$. A legal move reduces exactly one heap: choose $i$ with $x_{i}>0$ and replace $x_{i}$ by some $x_{i}^{\prime}<x_{i}$.
3.3 Sprague–Grundy theory and nim-sum
The Sprague–Grundy theorem states that every finite impartial game position is equivalent (under disjunctive sum) to a NIM heap of some size (its nimber/Grundy value). For NIM itself, the Grundy value of a position is the bitwise XOR (nim-sum) of its heap sizes.
Definition 3.3 (Nim-sum).
For heap sizes $x_{1},\dots,x_{h}$, define
$\operatorname{nim}(x_{1},\dots,x_{h})\;=\;x_{1}\oplus x_{2}\oplus\cdots\oplus x_{h},$
where $\oplus$ denotes bitwise XOR.
Theorem 3.4 (Classical NIM characterisation).
A NIM position $(x_{1},\dots,x_{h})$ is losing for the player to move iff $\operatorname{nim}(x_{1},\dots,x_{h})=0$. Moreover, from any winning position (nim-sum $\neq 0$) there is a move to a losing position (nim-sum =0=0).
4 Main results: impossibility and possibility
We separate the negative results for single-frame representations from the positive results obtained by adding recent history.
4.1 Mastery notions
We distinguish between two goals:
Definition 4.1 (Strong vs. weak mastery).
An agent achieves strong mastery of NIM if it plays optimally from every reachable position (equivalently, implements a winning strategy whenever one exists). It achieves weak mastery relative to an initial position $p_{0}$ if it plays optimally on the game tree reachable from $p_{0}$.
4.2 Single-frame impossibility results
In a single-frame representation, the network sees only the current position. To play optimally, it must in effect determine whether the nim-sum is zero and, if not, find a move producing nim-sum zero. The core obstacle is that global nim-sum computation requires parity-like behaviour across many bits, and parity is not in $\mathrm{AC}^{0}$ [9, 11, 19, 13].
Theorem 4.2 (Single-frame strong mastery requires parity; impossible in $\mathrm{AC}^{0}$).
Fix any encoding of NIM positions into Boolean strings of length $n$ such that each heap size is represented in binary with $\Theta(\log m)$ bits (so XOR of heap sizes corresponds to a parity-type computation over $\Theta(n)$ bits). No family of fixed-precision $\mathsf{NN}/\mathsf{RNN}/\mathsf{LTST}$ models (Definition 2.2) using only a single-frame input can achieve strong mastery of NIM for unbounded input sizes.
Proof sketch.
By Theorem 2.2, any such model can be simulated by $\mathrm{AC}^{0}$. Strong mastery implies that from every position the agent can decide (implicitly, via action choice) whether the position is winning or losing, which in NIM is equivalent to whether the global XOR (nim-sum) is zero (Theorem 3.4). This yields an $\mathrm{AC}^{0}$ computation of a parity-type predicate on $\Theta(n)$ bits, contradicting $\mathrm{PARITY}\notin\mathrm{AC}^{0}$ [9, 11, 19, 13]. ∎
Theorem 4.3 (Polynomial-time bounded search does not fix single-frame $\mathrm{AC}^{0}$ limitations).
Consider an AlphaZero-style agent that, at each node, uses fixed-precision single-frame $\mathrm{AC}^{0}$-constrained policy/value networks and performs any polynomial-time bounded tree search (including MCTS variants) whose node expansions and back-ups are polynomially bounded in the input length. Then such an agent still cannot achieve strong mastery of NIM for unbounded input sizes.
Proof sketch.
Intuitively, MCTS acts as a breadth-increaser: it allocates its simulation budget according to a policy prior and/or a learned value signal. If these signals cannot reliably distinguish nim-sum parity, then (in the regime of large heaps) the search remains essentially blind to the winning manifold: it has no robust criterion for concentrating visits on the unique restoring reductions rather than on the exponentially many non-restoring moves.
If the evaluation/policy signals are $\mathrm{AC}^{0}$-bounded and cannot reliably encode nim-sum parity, then the search lacks a systematic advantage in prioritizing winning moves as the branching factor grows. In NIM with large heaps (binary encoding), the fraction of winning moves can become extremely small in worst cases, and without parity information the search cannot consistently prune or focus to guarantee optimality within polynomial effort. A full formalization can be phrased as a reduction: any such search procedure that succeeds would induce (via transcript-to-circuit transformation) a polynomial-size constant-depth circuit family deciding nim-sum parity, again contradicting $\mathrm{PARITY}\notin\mathrm{AC}^{0}$. ∎
4.2.1 Implications for weak mastery
The above impossibility is for strong mastery. Weak mastery can sometimes be achieved from special initial positions with strong symmetry. For example, positions consisting of paired equal heaps permit mirroring strategies (Appendix D).
4.3 Multi-frame possibility results
The fundamental idea is to avoid computing the global nim-sum from scratch. Instead, we exploit that each move changes only one heap, so the nimber difference between consecutive positions is locally computable.
Definition 4.4 (Nimber difference).
Given consecutive positions $P$ and $P^{\prime}$ in a play sequence, define the nimber difference
$\Delta(P,P^{\prime})\;=\;\operatorname{nim}(P)\ \oplus\ \operatorname{nim}(P^{\prime}).$
In NIM, if $P^{\prime}$ is obtained from $P$ by changing heap $i$ from $x_{i}$ to $x_{i}^{\prime}$, then
$\Delta(P,P^{\prime})\;=\;x_{i}\oplus x_{i}^{\prime},$
a purely local computation on the changed heap.
Lemma 4.5 (Computing nimber difference in $\mathrm{AC}^{0}$ for bounded changes).
Fix a constant $t\in\mathbb{N}$. Let $P$ and $P^{\prime}$ be NIM positions whose heap vectors differ in at most $t$ heaps. Then $\Delta(P,P^{\prime})$ can be computed by an $\mathrm{AC}^{0}$ circuit of polynomial size (in the input length).
Proof.
See Appendix E. ∎
4.4 Two-frame local restoration rule
The next observation is that if, after our move, we make the opponent face a zero nim-sum position, then the opponent’s next move produces some nimber difference $\delta$. If we respond with a move that produces the same nimber difference $\delta$, we restore nim-sum zero.
Proposition 4.6 (Local restoration principle).
Suppose a player moves from a position $P$ with $\operatorname{nim}(P)=0$ to a position $Q$ (opponent’s turn). Let the opponent move to $Q^{\prime}$. Then
$\operatorname{nim}(Q^{\prime})\;=\;\Delta(Q,Q^{\prime}).$
If the player can respond with a move from $Q^{\prime}$ to $R$ satisfying
$\Delta(Q^{\prime},R)\;=\;\Delta(Q,Q^{\prime}),$
then $\operatorname{nim}(R)=0$.
Proof.
We use $\Delta(A,B)=\operatorname{nim}(A)\oplus\operatorname{nim}(B)$ and XOR associativity:
$\operatorname{nim}(R)\;=\;\operatorname{nim}(Q^{\prime})\oplus\Delta(Q^{\prime},R)\;=\;\Delta(Q,Q^{\prime})\oplus\Delta(Q,Q^{\prime})\;=\;0.$
Proposition 4.6 is structural: it describes a response pattern that keeps returning the game to nim-sum zero, provided the player can (i) compute $\delta=\Delta(Q,Q^{\prime})$ from the last two frames and (ii) choose a legal move $Q^{\prime}\to R$ that realizes the same $\delta$. In NIM, when the opponent changes exactly one heap, $\delta$ is simply $x_{i}\oplus x_{i}^{\prime}$, and testing whether a candidate response move realizes $\delta$ is a local check on the heap being changed.
The significance for learnability is representational. A two-frame policy network can condition its response on $(Q,Q^{\prime})$ and hence on the locally computable $\delta$, while a single-frame policy must reconstruct the relevant global information from $Q^{\prime}$ alone. In the experimental sections we use standard Monte Carlo tree search and ordinary simulated games; the multi-frame advantage comes from what the policy/value networks can infer from the input, rather than from imposing special constraints on simulation trajectories.
4.5 Implementation overview
An $\mathrm{AC}^{0}$ implementation of the two-frame delta and restoration components can be organised into:
  • a binary encoding scheme for heap vectors (Appendix C),
  • a constant-depth detection of which heaps changed between frames,
  • an $\mathrm{AC}^{0}$ circuit computing $\Delta$ on those changed heaps (Lemma 4.5),
  • an $\mathrm{AC}^{0}$ selection mechanism for a response move realizing a target $\delta$.
5 Learning impartial games with neural networks
Our results isolate a sharp distinction:
  • Weak learning: in special structured cases, single-frame agents can succeed via simple symmetries.
  • Strong learning: for general positions, strong mastery requires exploiting history and local invariants (nimber differences) rather than learning global parity.
5.1 Practical considerations
  • What single-frame networks miss. For large boards, fixed-precision bounded-depth networks cannot reliably encode the global nim-sum parity needed for strong mastery (Section 4.2). A single-frame policy/value network therefore has no direct way to base its decisions on the true win/loss structure of arbitrary positions.
  • What two frames add. With two consecutive positions $(Q,Q^{\prime})$, the agent can infer which heap changed and the induced nimber difference $\delta=\Delta(Q,Q^{\prime})$ (Lemma 4.5). Proposition 4.6 shows that conditioning the response move on this locally computable $\delta$ is enough to restore the zero nim-sum invariant after the opponent’s move.
  • Training signals and empirical protocols. Section 6 studies two complementary empirical settings. In Section 6.2 we isolate restoration as a supervised learning task using teacher–random transitions and train policies directly by cross-entropy to the set of optimal restoration moves. We also describe an AlphaZero-style teacher-guided MCTS training protocol (Section 6.3), which uses standard visit-count policy targets from MCTS and value targets from game outcomes. In all cases, simulations and rollouts are ordinary legal play; no special constraints are imposed on rollout moves beyond legality.
  • Coverage and robustness. To avoid collapsing onto a narrow set of deterministic optimal trajectories and to expose the agent to a broad range of transition patterns, we use a near-perfect teacher that plays optimally with high probability but makes a random legal move with small probability $\varepsilon$ (Section 5.2).
5.2 Teacher-guided training as a controlled source of transition data
Let $\pi^{*}$ denote an optimal NIM policy (equivalently, a policy that moves to a zero nim-sum position whenever possible). For a user-specified noise parameter $\varepsilon\in[0,1]$, we define a noisy expert (teacher) policy $\pi_{T}^{\varepsilon}$ as follows: on each move, with probability $1-\varepsilon$ the teacher selects an action according to $\pi^{*}$, and with probability $\varepsilon$ it selects a uniformly random legal move. During training the learner plays against $\pi_{T}^{\varepsilon}$ (and, when desired, uses the same teacher policy inside rollouts).
The occasional random moves serve two roles. First, they increase state-space coverage and prevent the training distribution from collapsing to a small set of optimal trajectories. Second, they generate a richer set of observed transitions $(Q,Q^{\prime})$, which is precisely the information exploited by two-frame representations and the local restoration principle (Proposition 4.6).
5.3 Approximate majority, parity, and why $\mathrm{AC}^{0}$ is the right “feasible” model
A key modelling issue is the following.
The apparent objection.
One could argue that neural networks can compute majority “in practice”, and therefore perhaps $\mathrm{TC}^{0}$-power (which includes majority gates) should be considered feasible. But then $\mathrm{TC}^{0}$ contains $\mathrm{AC}^{0}$, and yet parity remains empirically hard to learn and generalise [20, 21, 6]. How can majority be feasible while parity is not?
Resolution: feasible $\neq$ exact; sharpness matters.
The practical majority phenomenon is best understood as approximate majority: a classifier that gets the label right with high probability under natural noise/typical data distributions, but does not implement a perfectly sharp threshold at exactly $n/2$ for arbitrarily large $n$. Fixed precision and noisy training/inference make perfectly sharp thresholds unstable: tiny perturbations in weights, inputs, rounding, or intermediate activations can flip outcomes near the decision boundary. Thus, what is feasible is a robust rule: output $1$ when the fraction of $1$’s is clearly above $1/2$ and output 0 when clearly below, with an “uncertain band” around $1/2$.
$\mathrm{AC}^{0}$ already captures approximate majority.
It is important that $\mathrm{AC}^{0}$ can compute approximate majority-type predicates of this form (constant-depth polynomial-size approximate counting/thresholding), while it cannot compute parity (even with nontrivial correlation) [9, 11, 19, 13]. In other words, the practical learnability of majority-like behaviour does not force us to adopt $\mathrm{TC}^{0}$ as the feasible model.
Why majority does not “bootstrap” into parity in practice.
In $\mathrm{TC}^{0}$, one can compose exact threshold/majority gates to compute functions far beyond $\mathrm{AC}^{0}$ (including parity). But this relies on reliability: if each majority gate is only an approximation with even a small error probability, the error typically amplifies under composition, and parity is particularly unforgiving. Parity is maximally sensitive: flipping any single bit flips the output, so small independent errors in intermediate computations destroy correlation with the true parity. Put differently, parity requires a level of sharpness and global consistency that approximate majority does not provide.
Consequences for NIM learnability.
Optimal NIM play is governed by nim-sum, which is global XOR/parity structure (Theorem 3.4). Approximate majority-style computations do not help: even a small error rate in the parity bits of nim-sum produces systematically wrong move choices. Hence it is consistent that:
  • majority-like tasks are learnable/feasible, because coarse thresholds are robust and $\mathrm{AC}^{0}$-compatible;
  • parity-like tasks (and nim-sum) remain infeasible for practical fixed-precision bounded-depth networks;
  • multi-frame representations can restore feasibility by reducing the required parity computations to local parity on a constant number of changed heaps (Lemma 4.5).
This is precisely why an $\mathrm{AC}^{0}$-based model is a useful theoretical abstraction for practical learnability.
5.4 The Multi-Frame Advantage: From Global to Local Parity
While previous architectures, such as the original AlphaGo, utilized multi-frame stacks to capture temporal dynamics [16], we demonstrate that in impartial games, the transition from single-frame to multi-frame representation represents a fundamental shift in required circuit complexity.
  • Single-Frame Obstruction: A network restricted to a single position $P$ must compute the global nim-sum $\bigoplus_{i=1}^{n}x_{i}$. We prove this is equivalent to a PARITY-type computation, which is non-computable in $\mathrm{AC}^{0}$ [9].
  • Multi-Frame Discovery: By providing two frames $(P,P^{\prime})$, the network can isolate the changed heap $i$ and compute the local difference $\Delta=x_{i}\oplus x_{i}^{\prime}$. This local operation is provably within $\mathrm{AC}^{0}$ bounds (Lemma 4.5).
  • Empirical question. Section 6.2 tests whether two-frame policies learn the restoration response substantially more easily than single-frame policies when training data is organised around $(P\text{-position},\text{perturbed})$ transitions. We also discuss teacher-guided MCTS training (Section 6.3) as a more end-to-end but harder-to-bootstrap regime.
6 Experiments
We evaluate the multi-frame restoration hypothesis in two complementary settings. First, we construct a supervised restoration dataset from teacher–random transitions (Section 6.2), which directly labels the restoration response. Second, we describe an AlphaZero-style teacher-guided training regime using MCTS against a noisy expert (Section 6.3). Unless otherwise stated, both one-frame (1F) and two-frame (2F) agents share the same action parameterization and network architecture (Section 6.1).
6.1 Neural architecture and training objective
6.1.1 State and action representation
We use a fixed-width binary encoding of heap sizes with $k$ bits per heap. The single-frame input is the encoding of the current heap vector. The two-frame input concatenates the encodings of the previous and current heap vectors. When a previous frame is unavailable (e.g. on the first move of a game), we set the previous frame equal to the current position so that the input dimension is constant and the 2F policy remains well-defined.
The action space is parameterised as “set heap $i$ to value vv” for $i\in\{1,\dots,N\}$ and $v\in\{0,1,\dots,2^{k}-1\}$, giving $N\cdot 2^{k}$ action logits. Illegal actions (those with $v\geq x_{i}$) are masked before sampling or taking an argmax.
6.1.2 Network architecture
Both agents use the same MLP template with two heads: a policy head producing $N\cdot 2^{k}$ logits and a value head producing a scalar in [1,1][-1,1]. The two-frame network input has dimension $2Nk$; the single-frame network input has dimension $Nk$.
Component Type Dimensions Activation/Normalisation
Input State representation Two-frame: $2Nk$ bits (1F: $Nk$) None
Hidden 1 Fully Connected $2Nk\to H$ LayerNorm + ReLU
Hidden 2 Fully Connected $H\to H$ ReLU
Policy Head Linear Output $H\to N\cdot 2^{k}$ Log-Softmax
Value Head Linear Output $H\to 1$ Tanh
Table 1: Neural network architecture used in the experiments. $H$ denotes the hidden width (a fixed hyperparameter).
6.1.3 Training targets and loss
For each training state (single-frame) or state-pair (two-frame), MCTS produces a root visit-count distribution $\pi_{\mathrm{MCTS}}$. The policy head is trained to match this distribution (cross-entropy loss). The value head is trained to predict the final game outcome from the perspective of the player to move (mean squared error on targets $\pm 1$). The overall loss is the sum of policy and value losses (optionally with standard regularization).
6.1.4 Definition of mastery
We report two complementary notions of performance. Game-level performance is measured by win rate (and Elo) under the evaluation protocols of Section 6. Move-level performance is measured by functional correctness: on a test position with nonzero nim-sum, a move is counted as correct if it produces a successor state with nim-sum zero. Positions with zero nim-sum have no winning move (Theorem 3.4); when reporting functional correctness we therefore condition on nonzero nim-sum positions.
6.2 Supervised restoration learning from teacher–random transitions
This experiment isolates the restoration response as a supervised learning problem. In normal-play impartial games, a restoration response is the move that returns play to a losing ($P$-)position after the opponent departs from it. In Nim, this is equivalent to selecting an optimal (winning) move from the post-opponent position: if the previous position had nim-sum 0, then after any opponent move the position is winning, and any optimal reply restores nim-sum 0.
The goal is to test the core theoretical prediction of Sections 4.24.3 in a controlled, low-noise setting: does access to the transition $(p,p^{\prime})$ (two frames) make restoration learnable, even when one-frame models fail? This benchmark builds on the observation that parity-type invariants pose a challenge for reinforcement learning in impartial games [24, 25], while explicitly providing the transition information needed by Proposition 4.6.
6.2.1 Game and representations
We consider Nim with $N=20$ heaps. Each heap size is represented with $k=4$ bits (little-endian), so heap values lie in $\{0,\ldots,15\}$. Moves are represented in a fixed action space of size $N\cdot 2^{k}=20\cdot 16=320$: an action chooses a heap index $i$ and a new heap value $v$ (masked to legal moves $0\leq v<h_{i}$).
A 1-frame (1F) policy receives only the current position $p^{\prime}$ (80-bit input), while a 2-frame (2F) policy receives the concatenation $(p,p^{\prime})$ of the previous and current positions (160-bit input). Both agents use the shared MLP architecture described in Section 6.1 (two hidden layers of width $H=256$).
6.2.2 Dataset generation (teacher–random–teacher)
We generate a dataset of size $D=10^{6}$ examples of the form
$(p,p^{\prime})\mapsto\pi^{*}(\cdot\mid p^{\prime}),$
where $\pi^{*}(\cdot\mid p^{\prime})$ is a soft target distribution over optimal moves from $p^{\prime}$.
Each example is generated as follows:
  1. 1.
    Sample a random winning Nim position $s$ (nim-sum $\neq 0$).
  2. 2.
    Apply an optimal teacher move from $s$ to obtain $p$. For optimal play in Nim, $p$ has nim-sum 0 (a $P$-position for the next player).
  3. 3.
    Apply a uniformly random legal move from $p$ to obtain $p^{\prime}$.
  4. 4.
    Define the label $\pi^{*}(\cdot\mid p^{\prime})$ as the uniform distribution over all optimal (winning) moves from $p^{\prime}$.
This construction ensures that each training input $(p,p^{\prime})$ corresponds to the canonical restoration setting: $p$ is a $P$-position, the random move produces a winning position $p^{\prime}$, and the label is the set of restoration moves back to a $P$-position.
Compared to earlier empirical studies of impartial games [24, 25], we “help” the learner in two ways: (i) we represent heap sizes in binary rather than as raw integers, and (ii) we train on perfectly labelled restoration targets (the full set of optimal moves), rather than requiring the agent to infer these targets from sparse win/loss outcomes.
We split the dataset into 80% train / 10% validation / 10% test.
6.2.3 Training and checkpoints
We train each model with supervised learning using cross-entropy to the soft target distribution. We use Adam with learning rate $10^{-3}$ and batch size 512 for 200 epochs. To study performance as a function of training, we save $C=10$ checkpoints per model at epochs $\{1,23,45,67,89,112,134,156,178,200\}$.
6.2.4 Evaluation metrics
We evaluate in three complementary ways:
Restoration accuracy (validation/test).
Given a held-out example, we decode the model’s greedy action (argmax over legal moves). We count a prediction as correct if the chosen action lies in the support of $\pi^{*}(\cdot\mid p^{\prime})$ (i.e., it is one of the optimal restoration moves). This accounts for the fact that multiple optimal moves may exist.
Gameplay strength (checkpoint tournament, greedy play).
We form a pool of 20 agents (10 checkpoints from 1F and 10 from 2F) and run a round-robin tournament. Each pair plays 10 games (total 1900 games). Games start from a uniformly random nonterminal Nim position; the starting player is randomized. Agents play greedily from their policy logits (no search). We compute Elo ratings with initial rating 1500 for each agent and update factor $K=24$.
Gameplay strength with search (MCTS-per-move tournament).
To probe how restoration knowledge interacts with search, we also evaluate the same checkpoint pool under per-move Monte Carlo tree search. On every move, the acting agent runs a fixed-budget PUCT search (200 simulations per move; exploration constant $c_{\mathrm{puct}}=1.5$) using its policy network as the prior over legal actions. Leaf values are obtained by a rollout to termination using greedy play from the two checkpoint policies (one for each side); the rollout return is backed up with standard negation across plies. We run a head-to-head cross-group tournament (10 1F checkpoints $\times$ 10 2F checkpoints, 2 games per pair; 200 games total) and compute Elo as above.
6.2.5 Results
One-frame does not learn restoration; two-frame eventually does.
Figure 1 shows validation restoration accuracy as a function of epoch. The 1F model remains near chance throughout training, stabilizing around 0.079 (min 0.075, max 0.081 across checkpoints on the test set). In contrast, the 2F model learns the restoration mapping and reaches essentially perfect accuracy with enough training. On the test set (Figure 2), the 2F checkpoints achieve 0.209 at epoch 1, 0.538 at epoch 23, 0.876 at epoch 45, 0.989 at epoch 67, and 0.99998 by epoch 178 (remaining $\geq 0.99993$ at epoch 200).
Figure 1: Validation restoration accuracy vs. epoch on the restoration dataset ($N=20$, $k=4$, $D=10^{6}$). The 1F model stays near chance, while the 2F model exhibits a characteristic two-stage learning curve with an intermediate plateau around 0.54 before reaching near-perfect accuracy.
Figure 2: Test restoration accuracy of checkpointed agents ($N=20$, $k=4$, $D=10^{6}$). Checkpoints are saved at epochs $\{1,23,45,67,89,112,134,156,178,200\}$.
A learnability “plateau” explained by approximate-majority shortcuts.
The 2F learning curve is not monotone: it first rises to an intermediate plateau near $0.54$ and remains there for many epochs before “breaking through” to perfect restoration. This plateau is consistent with an approximate-majority shortcut on the least significant parity bit: in our generated data, the least significant bit of the post-random nim-sum is slightly imbalanced (empirically $P[\mathrm{LSB}(\mathrm{nim\_sum}(p^{\prime}))=1]\approx 0.5355$), so a policy that predicts this bit as a constant can achieve accuracy around 0.54 while still learning the easier aspects of the restoration response (e.g. which heap to modify and the higher bits of the target value). Only later does training force the model to represent the full parity structure needed for exact restoration, at which point accuracy quickly approaches 1. This behaviour reflects the broader theme of Section 5.3: approximate majority-type correlations are $\mathrm{AC}^{0}$-compatible and easy to pick up, while exact parity requires sharper global consistency.
Greedy play: two frames yield a modest Elo advantage.
In the greedy round-robin tournament (no search), 2F agents win 59.6% of cross-group games (596/1000; 95% CI 56.5%–62.6%). The mean Elo of 2F checkpoints is 1562.5 versus 1437.5 for 1F (a +125.1 Elo gap).
Search amplifies the two-frame advantage.
Under per-move MCTS with a fixed search budget (200 simulations per move), the advantage of 2F priors becomes much larger. 2F wins 79.5% of cross-group games (159/200; 95% CI 73.4%–84.5%), and the mean Elo gap increases to +175.0 (1587.5 for 2F vs. 1412.5 for 1F). Figure 3 shows the resulting Elo ordering across checkpoints.
Figure 3: Head-to-head Elo of checkpointed agents under MCTS-per-move evaluation ($N=20$, $k=4$). Each move is selected by PUCT with 200 simulations and rollouts using greedy policy play from the two checkpoints.
Overall, these results support the theoretical narrative: single-frame models fail to learn the parity-governed optimal response even under perfect supervision, while two-frame models can learn the restoration rule from transition data. Moreover, once a reliable restoration prior is learned, combining it with bounded search substantially increases effective playing strength.
6.3 Teacher-guided MCTS training against a noisy teacher
This subsection describes an AlphaZero-style training protocol in which the learner plays against a near-perfect noisy teacher and selects actions via MCTS. We include it as a controlled end-to-end baseline. In practice, bootstrapping strong play from terminal win/loss outcomes can be challenging even with a teacher; the supervised restoration benchmark of Section 6.2 provides a cleaner test of the multi-frame restoration hypothesis.
This section describes an experimental protocol designed to test the theoretical prediction of Sections 4.24.3: providing a short history window should make NIM substantially easier to learn for fixed-precision bounded-depth policy/value networks.
6.3.1 Game family and start-state distribution
We consider the family of NIM games with $N$ heaps and heap sizes in $\{0,1,\dots,2^{k}-1\}$. At the start of each training episode we sample a nonterminal position by drawing each heap size independently and uniformly from $\{0,1,\dots,2^{k}-1\}$ and resampling if all heaps are zero. For evaluation we optionally stratify results by whether the initial nim-sum is zero or nonzero (Theorem 3.4).
6.3.2 Near-perfect teacher
We use the noisy expert policy $\pi_{T}^{\varepsilon}$ from Section 5.2. When $\varepsilon=0$ the teacher plays perfectly; for $\varepsilon>0$ the teacher occasionally makes random moves, which increases the diversity of visited states and transitions.
6.3.3 Training procedure: MCTS against the teacher
We train two agents independently: a single-frame agent whose networks take the current position as input, and a two-frame agent whose networks take the previous and current positions as input. Both agents are trained in an AlphaZero-style loop:
  1. (1)
    Play episodes against the teacher. On the agent’s turn, select actions using Monte Carlo tree search (MCTS) with a fixed simulation budget.
  2. (2)
    Use the root visit-count distribution produced by MCTS as the policy training target for the corresponding state (or state-pair).
  3. (3)
[…]
  4. (4)
    Update network parameters by stochastic gradient descent on mini-batches from the replay buffer.
Rollouts and back-ups.
Within each MCTS simulation, leaf evaluation can be obtained either from the current value network or from a rollout that plays the game forward to termination using a simple default policy (e.g. the current policy network with added exploration and/or the teacher for the opponent). In all cases, the return of the simulation is backed up along the path in the standard way. No additional constraints are imposed on rollout moves beyond legality.
6.3.4 Evaluation metrics
We report:
  • Win rate vs. a perfect teacher ($\varepsilon=0$), which measures performance against optimal play.
  • Win rate vs. the noisy teacher ($\varepsilon$ matching training), which measures robustness and exploitation of occasional mistakes.
  • Head-to-head win rate and Elo difference between the 1F and 2F agents under the same search budget and alternating first player.
  • Move optimality accuracy on random test positions with nonzero nim-sum: the fraction of states where the agent’s chosen move produces nim-sum zero.
For Elo reporting we use the standard logistic model. Given an observed win rate $w$ for the two-frame agent in a head-to-head match, the implied Elo difference is
$\Delta_{\mathrm{Elo}}\,=\,-400\log_{10}\!\Big(\frac{1}{w}-1\Big),$
with $w$ clipped away from 0 and $1$ for numerical stability.
7 Discussion and broader implications
Our main message is that multi-frame representations can turn globally parity-governed decision rules into locally checkable transition rules, making them feasible for fixed-precision bounded-depth neural inference. We briefly discuss two directions where this perspective may be useful beyond the specific NIM setting.
7.1 Multi-frame approaches to mathematical discovery
The $\mathrm{AC}^{0}$ perspective is not limited to game playing. It provides a general method for diagnosing when a direct neural approach is blocked by parity-like constraints and for designing alternative representations that bypass the barrier.
This principle guided our work on discovering large Condorcet domains [12, 23, 22]. The long-standing record domain (Fishburn) involves parity-like constraints across many triples; single-frame neural search struggled even at small parameters. Multi-frame and locally checkable invariants transformed the search from an $\mathrm{AC}^{0}$-infeasible global parity constraint into tractable local constraints that guide exploration.
7.2 Beyond NIM: other impartial games
By Sprague–Grundy theory, every finite impartial game position has a nimber. However, the computational complexity of determining nimbers varies widely. Some impartial games are PSPACE-complete (e.g. Geography) [14]. In such cases, even polynomial-time algorithms (let alone $\mathrm{AC}^{0}$-bounded inference) cannot compute optimal play from arbitrary positions.
Our message is not that $\mathrm{AC}^{0}$ is universally sufficient, but that it is a useful model for diagnosing when learning failure is due to parity-like global constraints and for motivating representation/search strategies that reduce global constraints to local checks.
8 Conclusion
We proposed a circuit-complexity model for practical neural inference: constant-depth, polynomial-size networks with fixed-precision parameters. We showed that these models are simulable by $\mathrm{AC}^{0}$, which makes global parity-type computations (and hence NIM’s nim-sum) infeasible in the strong, worst-case sense.
This yields a clean explanation for persistent learning failures under single-frame training: strong mastery of NIM requires reliable access to global nim-sum information, which fixed-precision bounded-depth networks cannot represent, and polynomial-time bounded search does not remove this obstruction. On the positive side, we proved that a short history window exposes locally computable nimber differences, and that a player can restore the zero nim-sum invariant by matching the observed nimber difference between consecutive frames.
Our experiments support the restoration viewpoint in a controlled setting. In a supervised teacher–random restoration dataset for 20-heap, 4-bit NIM, the one-frame policy remains near chance, while the two-frame policy learns restoration via a characteristic two-stage curve (an intermediate plateau around 0.54 followed by near-perfect accuracy) and achieves a clear head-to-head advantage over the one-frame baseline, especially when combined with bounded per-move MCTS (Section 6.2). At the same time, more end-to-end teacher-guided MCTS training remains a challenging regime to bootstrap from terminal win/loss signals, highlighting the value of intermediate, transition-aligned learning signals for restoration behaviour.
More broadly, the $\mathrm{AC}^{0}$ perspective offers a diagnostic tool: when learning fails, the bottleneck may be a parity-like global constraint rather than a lack of data or optimization. Multi-frame inputs and locally checkable invariants provide a principled route to bypass such barriers in other combinatorial domains.
Appendix A Formal definitions of fixed-precision neural models
This appendix provides formal definitions of the fixed-precision neural models used throughout the paper.
A.1 Preliminaries: fixed precision and representable parameters
Fix constants $W,D\in\mathbb{N}$. Define the fixed finite set of representable parameters
$\mathcal{Q}_{W,D}=\Big\{\frac{k}{D}:k\in\{0,1,\dots,W\}\Big\}.$
In particular, the set of allowable weights/thresholds does not depend on input length $n$. This is the mathematical formalization of “32-bit precision” (or any constant word size): there are only finitely many representable constants.
We will allow negations on wires (equivalently, $\mathrm{AC}^{0}$ has NOT gates). This is standard in circuit complexity and avoids conflating two independent issues: (i) parity hardness vs (ii) monotonicity restrictions.
A.2 Feedforward model
Definition A.1 (Fixed-precision constant-depth neural network).
A fixed-precision $\mathsf{NN}$ of depth $L$ on $n$ Boolean inputs is specified by:
  • Layer widths $m_{0}=n$, $m_{1},\dots,m_{L}$, where $m_{\ell}\leq\mathrm{poly}(n)$.
  • For each layer $\ell\in\{1,\dots,L\}$ and unit $i\in\{1,\dots,m_{\ell}\}$:
    • weights $w_{\ell,i,j}\in\mathcal{Q}_{W,D}$ for $j\in\{1,\dots,m_{\ell-1}\}$,
    • a threshold $\theta_{\ell,i}\in\mathcal{Q}_{W,D}$,
    • an optional negation bit $\eta_{\ell,i,j}\in\{0,1\}$ indicating whether input $j$ is negated before weighting.
The computation proceeds as follows. Let $y^{(0)}=x\in\{0,1\}^{n}$. For $\ell\geq 1$,
$y^{(\ell)}_{i}\;=\;\mathbf{H}\!\Big(\sum_{j=1}^{m_{\ell-1}}w_{\ell,i,j}\cdot(y^{(\ell-1)}_{j}\oplus\eta_{\ell,i,j})\;-\;\theta_{\ell,i}\Big),$
where $\mathbf{H}(z)=1$ if $z\geq 0$ and $\mathbf{H}(z)=0$ otherwise. The network output is $y^{(L)}$ (or a designated subset of its coordinates).
A.3 Recurrent model (finite window)
Definition A.2 (Fixed-precision recurrent neural network with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-precision $\mathsf{RNN}$ consists of:
  • input bits $x_{t}\in\{0,1\}^{n}$ at time $t$,
  • a hidden state $h_{t}\in\{0,1\}^{m}$ for some $m\leq\mathrm{poly}(n)$,
  • an output $o_{t}\in\{0,1\}^{k}$ for $k\leq\mathrm{poly}(n)$,
  • fixed-precision parameters (weights/thresholds in $\mathcal{Q}_{W,D}$ and optional input negations) shared across time,
such that each update $h_{t+1}$ and $o_{t}$ is computed by a constant-depth fixed-precision threshold network over inputs drawn from $\{x_{t^{\prime}},h_{t^{\prime}}:t-T<t^{\prime}\leq t\}$. Equivalently, unrolling for $T$ steps yields a constant-depth feedforward computation with depth $O(L\cdot T)$.
A.4 Attention-style model (finite window)
Definition A.3 (Fixed-precision LTST model with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-precision $\mathsf{LTST}$ model processes a sequence $(x_{t-T+1},\dots,x_{t})$ by:
  • computing short-term and long-term attention scores using fixed-precision parameters in $\mathcal{Q}_{W,D}$,
  • combining attended summaries through a constant number of layers of fixed-precision threshold computations,
  • producing outputs $o_{t}\in\{0,1\}^{k}$.
We assume that all attention coefficients, thresholds, and intermediate gating parameters are drawn from $\mathcal{Q}_{W,D}$ and that the number of attention heads/layers is constant, while the total number of units is polynomial in $n$. Unrolling the fixed window $T$ yields a constant-depth feedforward threshold computation of depth $O(L\cdot T)$.
Remark A.4 (Why these formal definitions match $\mathrm{AC}^{0}$).
Theorem 2.2 applies because $L,T,W,D$ are constants and each unit’s parameters come from a finite set independent of $n$. Appendix B gives an explicit circuit construction.
Appendix B An $\mathrm{AC}^{0}$ construction for simulating fixed-precision networks
This appendix gives a clean circuit construction proving Theorem 2.2.
B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed precision
We first show that a single fixed-precision threshold unit (as in Definition A.1) is computed by a depth-2 $\mathrm{AC}^{0}$ circuit of polynomial size.
Lemma B.1 (Fixed-precision threshold gate $\in\mathrm{AC}^{0}$).
Fix constants $W,D\in\mathbb{N}$. Let $z_{1},\dots,z_{m}\in\{0,1\}$ be Boolean inputs, and let $w_{1},\dots,w_{m},\theta\in\mathcal{Q}_{W,D}$. Consider the Boolean function
$f(z_{1},\dots,z_{m})\;=\;\mathbf{H}\!\Big(\sum_{j=1}^{m}w_{j}z_{j}-\theta\Big).$
Then $f$ is computable by a depth-2 $\mathrm{AC}^{0}$ circuit of size $O(m^{K})$, where $K:=WD$ is a constant (depending only on $W,D$).
Proof.
Multiply the inequality by $D$ to clear denominators. Write $w_{j}=a_{j}/D$ and $\theta=b/D$ with $a_{j}\in\{0,1,\dots,W\}$ and $b\in\{0,1,\dots,W\}$. Then
$f(z)=1\quad\Longleftrightarrow\quad\sum_{j=1}^{m}a_{j}z_{j}\;\geq\;b.$
If $b=0$ the function is identically $1$. Assume $b\geq 1$.
Because each $a_{j}$ is a nonnegative integer and each active input contributes at least $1$ to the left-hand side whenever $a_{j}\geq 1$, any satisfying assignment contains a subset of at most $b$ active variables whose coefficients already sum to at least $b$. More formally: if $\sum_{j}a_{j}z_{j}\geq b$, let $S=\{j:z_{j}=1\text{ and }a_{j}\geq 1\}$. Choose any multiset-free subset $T\subseteq S$ by greedily adding indices until the partial sum reaches $\geq b$. Since each added index increases the sum by at least $1$, this process stops after at most $b\leq W$ steps, hence $|T|\leq b\leq W\leq WD=K$.
Therefore,
$f(z)=1\quad\Longleftrightarrow\quad\bigvee_{\begin{subarray}{c}T\subseteq[m]\\ |T|\leq K\\ \sum_{j\in T}a_{j}\geq b\end{subarray}}\ \bigwedge_{j\in T}z_{j}.$
This is a depth-2 DNF: an OR of ANDs. The number of terms is at most $\sum_{r=0}^{K}\binom{m}{r}=O(m^{K})$, polynomial since $K$ is constant. Unbounded fan-in AND/OR gates give an $\mathrm{AC}^{0}$ circuit of depth $2$ and size $O(m^{K})$. ∎
Remark B.2 (Negations).
If some inputs are negated before weighting, i.e. $(z_{j}\oplus\eta_{j})$, this simply replaces $z_{j}$ by either $z_{j}$ or $\neg z_{j}$ in each conjunction term. Since $\mathrm{AC}^{0}$ allows NOT gates, Lemma B.1 still applies.
B.2 From one unit to a full network
Lemma B.3 (One layer of fixed-precision units is in $\mathrm{AC}^{0}$).
Fix $W,D$ and let a layer consist of $m^{\prime}$ threshold units whose inputs are outputs of a previous layer of width $m$, with $m,m^{\prime}\leq\mathrm{poly}(n)$. Then the layer can be simulated by an $\mathrm{AC}^{0}$ circuit of constant depth and polynomial size.
Proof.
Apply Lemma B.1 independently to each unit. Each unit yields a depth-2 circuit of size $O(m^{K})$. Combine these $m^{\prime}$ circuits in parallel; the total size is $m^{\prime}\cdot O(m^{K})=\mathrm{poly}(n)$ and depth remains $2$. ∎
Proof of Theorem 2.2.
For a feedforward $\mathsf{NN}$ of constant depth $L$, repeatedly replace each layer by the $\mathrm{AC}^{0}$ circuit from Lemma B.3, wiring outputs of layer $\ell-1$ circuits as inputs to layer $\ell$ circuits. Depth multiplies by a constant factor: overall depth is $O(L)$ (constant), and size remains polynomial as a composition of polynomial-size layers.
For $\mathsf{RNN}$ and $\mathsf{LTST}$ with time window $T$, unroll the computation for $T$ steps (constant) to obtain a feedforward computation of depth $O(LT)$ (still constant). Apply the same layer-by-layer replacement. ∎
Appendix C Implementation details (overview)
This appendix summarizes the components needed to implement the two-frame delta computation and the local restoration response within $\mathrm{AC}^{0}$ constraints.
  • Encoding. Represent each heap size in binary using a fixed block of bits; represent a position as the concatenation of blocks.
  • Difference detection. Given two frames $P,P^{\prime}$, compute in parallel which heap blocks differ (bitwise XOR within each block, then OR across the block).
  • Compute $\Delta$. If positions differ in at most $t$ heaps, compute the XOR of the changed-heap values (Appendix E). Since $t$ is constant, this is constant-depth.
  • Move selection. Given $\delta$ and current position, in parallel for each heap compute the candidate new size $x_{i}^{\prime}=x_{i}\oplus\delta$ and test legality ($x_{i}^{\prime}<x_{i}$). Use a fixed tie-breaking rule to output one legal move (e.g. smallest index).
All subroutines are implementable in $\mathrm{AC}^{0}$ because comparisons and equality on $O(\log m)$-bit blocks are depth-2 with unbounded fan-in.
Appendix D Weak mastery examples
Example D.1 (Pairing strategy).
Consider positions $(a,a,b,b,c,c,\dots)$ consisting of pairs of equal heaps. The second player can mirror the first player’s move on the matching heap, maintaining the paired structure until termination, hence winning. This strategy is implementable by simple local equality checks and does not require global nim-sum computation.
Appendix E Proof of Lemma 4.5
We provide an explicit $\mathrm{AC}^{0}$ construction for computing nimber differences between positions differing in at most $t$ heaps.
Proof of Lemma 4.5.
Let positions be encoded as $h$ blocks of length $B$ bits (heap sizes in binary), so total input length is $n=hB$.
Step 1: identify changed heaps. For each heap $i$ and bit position $k\in[B]$, compute $d_{i,k}=x_{i,k}\oplus x^{\prime}_{i,k}$. Then $c_{i}=\bigvee_{k\in[B]}d_{i,k}$ indicates whether heap $i$ changed. This is depth-2 (XOR can be written as depth-2 using AND/OR/NOT on two bits; the OR over $B$ bits is unbounded fan-in).
Step 2: assume at most $t$ changes. By promise, $\sum_{i}c_{i}\leq t$ with $t$ constant.
Step 3: compute per-changed-heap XOR values. For each heap $i$, define the block XOR value $y_{i}=x_{i}\oplus x^{\prime}_{i}$ (bitwise XOR across the $B$ bits). Each output bit is XOR of two bits, hence $\mathrm{AC}^{0}$.
Step 4: XOR the at-most-$t$ nonzero blocks. Since at most $t$ heaps changed, $\Delta=\bigoplus_{i:c_{i}=1}y_{i}$ is XOR of at most $t$ $B$-bit numbers. For each bit position $k$, the $k$-th bit of $\Delta$ is parity of at most $t$ bits $\{y_{i,k}:c_{i}=1\}$. Because $t$ is constant, this parity can be computed by a constant-size DNF (or by hardwiring the truth table) and thus is in $\mathrm{AC}^{0}$.
The overall circuit has constant depth and polynomial size in $n$. ∎