### v3

# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

Abstract
[…]
Keywords: impartial games; NIM; AlphaZero; circuit complexity; quantised inference; fixed latency; planning; Monte Carlo tree search.
1 Introduction
Game play has historically served as the “Formula 1” of artificial intelligence: a high-speed, controlled proving ground where engineering limits are tested and new paradigms are born. From Deep Blue in chess to AlphaGo in Go, games have repeatedly driven advances in search, planning, and reinforcement learning [21, 23, 22, 20]. Negative results have been equally influential: Minsky and Papert’s impossibility result for single-layer perceptrons on parity/XOR redirected the field towards multilayer architectures [11].
History demonstrates that optimal algorithms are often reverse-engineered from such lower bounds: Nesterov’s accelerated method and worst-case optimal join algorithms both emerged from explicit complexity limits [15, 16]. We adopt this methodology by using circuit complexity not merely to predict failure, but to derive the minimal architectural priors required for scalable and robust reasoning. Our contribution is an explicit barrier $\to$ bypass methodology: each barrier we prove is paired with a targeted representational or hybrid step motivated by the barrier mechanism.
Today, despite the success of AlphaZero-style methods, such a barrier appears in impartial games such as NIM. Motivated by Harvey Friedman’s proposal of NIM as an AI challenge [4], several empirical studies asked whether standard AlphaZero pipelines can master this setting [31, 29, 32]. A recurring phenomenon is that, despite NIM’s simple rules, learning optimal play is blocked by a global parity invariant (the nim-sum).
This paper formalises that barrier using circuit complexity and derives the structural bypasses. The central observation is that while the state can be hard to process, the transition can be easy: augmenting with short history converts a global hardness obstacle into a local verification task. Figure 1 summarises this viewpoint for NIM.
Input $P_{t}$ (Global) Single-Head Network $\times$ Cannot compute
global parity
[…]
(a) Barrier: Single-Frame, Single-Head Input $(P_{t-1},P_{t})$ (Transition) Single-Head Network $\checkmark$ Computes Local
Difference $\Delta$
(b) Bypass 1: Multi-Frame Input Input $P_{t}$ (Global) Head $\pi_{0}$ Head $\pi_{k}$ $\vdots\checkmark$ Heads recover
$B$ parity bits
(c) Bypass 2: Multi-Head Output
[…]
1.1 Contributions
  • $\mathrm{AC}^{0}$ containment for fixed-scale quantised inference. We formalise three families of bounded-latency neural models (feedforward threshold networks, finite-window recurrent networks, and finite-window attention-style models) operating in a fixed-scale quantised regime (bounded range and fixed quantisation grid independent of input length), and show they are simulable by $\mathrm{AC}^{0}$ circuits (Theorem 2.4).
  • Single-frame impossibility for strong mastery. Under this model, no single-frame policy/value network can strongly master NIM in the worst case (Theorem 5.2).
  • Planning analysis and the single-policy bottleneck. We analyse MCTS-style rollouts as an interface between a weak evaluator and a long-horizon game. With $B$ distinct rollout heads one can recover all $B$ bits of the nim-sum by depth amplification (Proposition 5.9), whereas a single rollout head reveals at most one fixed linear functional of the nim-sum regardless of search budget (Proposition 5.10 and Theorem 5.14).
  • Two-frame bypass via local differences and restoration. With two consecutive frames, the agent can compute a local nimber difference $\Delta(P,P^{\prime})$ in $\mathrm{AC}^{0}$ (Lemma 5.4) and implement a restoration rule that maintains $\operatorname{nim}=0$ (Proposition 5.5), yielding a universal-verifier rollout policy (Proposition 5.13).
  • Empirical validation and learnability discussion. On 20-heap, 4-bit NIM with $10^{6}$ supervised restoration examples, a one-frame model stays near chance while a two-frame model learns restoration (Section 7.2). In a complementary single-frame experiment, multi-head FSM-controlled shootouts achieve perfect win/loss position classification while the single-head baseline remains bottlenecked by one-bit information (Section 7.3). We also discuss why $\mathrm{AC}^{0}$ is a useful abstraction for what is robustly feasible under fixed-scale quantised inference (Section 6.3).
1.2 Scope and limitations
Our theorems are worst-case representational statements about what can be computed under explicit architectural constraints. They are not training-complexity results, and they intentionally abstract away from several practical details.
  • Fixed-scale quantisation / fixed latency. The $\mathrm{AC}^{0}$ simulation (Theorem 2.4) assumes constant depth and a fixed-scale, bounded-range quantisation regime: weights/thresholds lie on a fixed grid whose range/resolution do not scale with the input length (cf. Remark 2.2). Architectures that rely on fan-in dependent re-normalisation, biases/thresholds growing with dimension, or depth growing with $n$ fall outside this abstraction.
  • Theoretical idealisation vs. practical normalisation layers. Our formal abstraction targets the logical reasoning layers that must compute parity-type invariants. The experimental networks also include LayerNorm and Softmax (implemented as Log-Softmax), which are statistical normalisation layers with global reductions and therefore exceed strict $\mathrm{AC}^{0}$ as literal circuit primitives. We treat these layers as calibration/normalisation operations, distinct from the logical reasoning layers, rather than as sources of parity-computing power. Empirically, this distinction is consistent with Section 7.2: the one-frame model still fails despite having the same normalisation layers.
  • Stylised planning interface. Our rollout results model planning as querying deterministic rollout outcomes under a specified policy head. This captures the information that a weak evaluator can expose through repeated simulation, but it does not model all details of practical MCTS (stochastic policies, exploration noise, learned value backup, etc.).
  • No guarantee of learnability. The multi-policy-head construction proves existence of a strong-mastery algorithm under multiple heads; we do not claim that self-play efficiently discovers the required structured heads. Our experiments therefore focus on a controlled supervised restoration signal that isolates the representational difference between one-frame and two-frame inputs.
1.3 Paper organisation
Section 2 introduces the fixed-scale quantised inference model and the $\mathrm{AC}^{0}$ simulation theorem. Section 3 situates our contribution in circuit complexity, modern sequence models, and reinforcement learning for combinatorial games. Section 4 recalls NIM and the nim-sum characterisation. Section 5 presents the main impossibility and bypass results, including the rollout analysis in Section 5.4. Section 6 discusses modelling choices and learnability considerations. Section 7 reports experiments. Appendices contain formal model definitions and proofs.
2 Theoretical framework and neural network models
2.1 Fixed-scale quantised inference and $\mathrm{AC}^{0}$ simulation
We consider neural inference models that are (i) bounded depth, (ii) polynomial size, and (iii) operate in a fixed-scale quantised regime: all weights and thresholds lie on a fixed quantisation grid of bounded dynamic range that does not change with the input length. This is a stylised abstraction of deployment-time low-bit inference (e.g. INT8/fixed-point) with fixed scaling factors, where one does not re-normalise weights by fan-in as the problem size grows.
Concretely, we fix constants:
  • a quantisation resolution parameter captured by a fixed denominator $D\in\mathbb{N}$ (grid step $1/D$),
  • a dynamic-range bound $W\in\mathbb{N}$ (maximum magnitude $W/D$),
  • and (for recurrent/attention models) a finite time window $T\in\mathbb{N}$.
We define
$\mathcal{Q}_{W,D}\;=\;\Big\{\frac{\ell}{D}\ :\ \ell\in\{0,1,2,\dots,W\}\Big\},$
a finite set independent of input length.
Each neuron computes a thresholded weighted sum of Boolean inputs. Formal definitions for $\mathsf{NN}$, $\mathsf{RNN}$ and $\mathsf{LTST}$ are in Appendix A.
Definition 2.1 (Neural networks with fixed-scale quantised parameters (informal)).
A fixed-scale quantised $\mathsf{NN}$ is a constant-depth feedforward network of polynomially many threshold neurons, whose weights and thresholds lie in $\mathcal{Q}_{W,D}$. Analogously, fixed-scale quantised $\mathsf{RNN}$ and $\mathsf{LTST}$ models have constant depth and polynomial size per time step and operate within a fixed time window $T$.
Why fixed-scale quantisation suggests an $\mathrm{AC}^{0}$ abstraction.
The key structural property is fixed scale: the smallest nonzero weight is $1/D$ and the largest threshold is $W/D$, with $W,D$ independent of the input length. Consequently, whenever a threshold unit outputs $1$, there is a satisfying witness consisting of at most $W$ input literals whose contributions already cross the threshold. This yields a constant-width DNF expansion for each unit (Lemma B.1), and composing a constant number of layers yields an $\mathrm{AC}^{0}$ simulation (Theorem 2.4).
We emphasise that this is not a statement about floating-point arithmetic per se. The abstraction intentionally excludes scale-adaptive regimes where thresholds/biases grow with fan-in or where weights shrink with dimension (e.g. $1/n$ or $1/\sqrt{n}$ normalisations), which move the analysis toward threshold-circuit classes such as $\mathrm{TC}^{0}$.
Remark 2.2 (Scope of the $\mathrm{AC}^{0}$ abstraction).
Theorem 2.4 formalises a fixed-latency fixed-scale quantised regime in which each unit uses parameters drawn from a bounded-range grid $\mathcal{Q}_{W,D}$ with $W,D$ independent of the input length. This yields the constant-witness property exploited in Lemma B.1.
The abstraction is deliberately conservative: it does not model architectures that rely on fan-in dependent re-scaling, numerically large signed biases/thresholds that grow with dimension, or deep computations whose depth scales with $n$. Allowing such scale-adaptive behaviour leads naturally to richer threshold-circuit models (e.g. $\mathrm{TC}^{0}$), in which parity-like computations can become representable.
Remark 2.3 (Local vs. global parity).
When we say “parity is hard” we mean global parity across the full Boolean input of length $n=NB$. In contrast, parity on $O(1)$ bits is trivial in $\mathrm{AC}^{0}$ (hardwire the truth table), and more generally XOR of a constant number of $B$-bit blocks is $\mathrm{AC}^{0}$-computable. This local/global separation is exactly what underlies the two-frame bypass via the nimber difference $\Delta(P,P^{\prime})$.
Property of deployed inference Captured here Not captured here
Fixed latency / bounded depth constant depth $L$ (and window $T$) depth growing with $n$
Fixed-scale quantisation (bounded range) weights/thresholds in fixed grid $\mathcal{Q}_{W,D}$ fan-in dependent scaling or thresholds growing with $n$
Parallel bit operations unbounded fan-in AND/OR/NOT exact global counting / exact parity on unbounded inputs
Signed effects via input negations (Lemma B.1) large signed dot-product accumulation
Notation convention (input length).
We write $N$ for the number of heaps and $B$ (or $k$ in experiments) for the bit-width used to encode each heap size. The resulting Boolean input length is $n:=NB$. When we discuss $\mathrm{AC}^{0}$ circuit size/depth, $n$ always refers to this Boolean input length.
Circuit complexity primer.
$\mathrm{AC}^{0}$ is the class of Boolean functions computable by constant-depth, polynomial-size circuits with unbounded fan-in AND/OR/NOT gates. A classical result is that $\mathrm{PARITY}\notin\mathrm{AC}^{0}$. In contrast, $\mathrm{TC}^{0}$ augments $\mathrm{AC}^{0}$ with unbounded fan-in threshold gates and can compute parity. Our focus here is a fixed-scale quantised regime in which thresholds do not scale with fan-in; in that regime, each threshold unit admits a constant-witness DNF expansion and the overall computation collapses to $\mathrm{AC}^{0}$ (Theorem 2.4).
2.2 Simulation by circuits
We state a single theorem that applies to all three models.
Theorem 2.4 (Simulation of fixed-scale quantised neural models by $\mathrm{AC}^{0}$).
Fix constants $W,D\in\mathbb{N}$ and a time window $T\in\mathbb{N}$. Let $\{\mathcal{M}_{n}\}_{n\geq 1}$ be any family of models of type $\mathsf{NN}$, $\mathsf{RNN}$ (with window $T$), or $\mathsf{LTST}$ (with window $T$) such that:
  1. (1)
    (Polynomial size) the number of neurons/units in $\mathcal{M}_{n}$ is at most $\mathrm{poly}(n)$.
  2. (2)
    (Constant depth) the number of layers is a fixed constant $L$ (independent of $n$); for $\mathsf{RNN}/\mathsf{LTST}$ the unrolled computation uses at most $T$ time steps.
  3. (3)
    (Fixed-scale quantisation) all weights and thresholds lie in $\mathcal{Q}_{W,D}$.
Then the Boolean function computed by $\mathcal{M}_{n}$ can be computed by an $\mathrm{AC}^{0}$ circuit family of polynomial size and constant depth (depending only on $L,T,W,D$).
3 Related Work
Circuit complexity and neural networks.
The correspondence between neural computation and circuit classes is foundational. Parity is outside $\mathrm{AC}^{0}$ [5, 7], while constant-depth threshold circuits ($\mathrm{TC}^{0}$) are more expressive. Recent analyses connect modern attention architectures to this lens [10, 6]. In particular, Merrill and Sabharwal show that parity-like behaviour remains difficult for standard transformer computation without additional structured reasoning channels [10]. Our contribution differs in focus: we model the control policy in RL in a fixed-latency, fixed-scale quantised inference regime and study the consequences for planning.
Quantized inference and deployment constraints.
Our fixed-scale assumption is aligned with practical low-bit deployment regimes. Integer-arithmetic-only inference with fixed scaling and bounded dynamic range is standard in efficient neural deployment pipelines [13, 14]. We use this as a conservative theoretical abstraction: not to model every floating-point detail, but to capture a regime where threshold scales do not grow with input dimension.
Reinforcement learning in combinatorial games.
While RL has mastered domains such as Go and chess, impartial games remain a useful stress test because optimality is governed by exact algebraic invariants rather than approximate pattern matching. Friedman [4] explicitly proposed NIM as an AI challenge for this reason. Zhou and Riis [31, 32] reported empirical failures of standard DQN/PPO-style pipelines on NIM. Our results provide a formal impossibility account that explains those observations in a fixed-scale quantised, fixed-latency regime.
History and memory in RL.
Using short history windows is standard in partially observable settings, where stacked frames recover hidden variables such as velocity [12]. Our point is different and perhaps surprising: even in Markovian, fully observable games—where the current state is informationally sufficient and the past is theoretically irrelevant—history can reduce computational complexity by exposing local differences that would otherwise require expensive search or unbounded-depth inference to extract from the current state alone. In our setting, two consecutive frames turn a parity-heavy global decision into an $\mathrm{AC}^{0}$-computable local restoration rule: the previous frame effectively donates, for free, the computational work that search would need to perform.
4 Impartial games and NIM
We briefly recall standard definitions; see [2] for extensive background.
4.1 Impartial games
Definition 4.1 (Impartial game).
An impartial game is a pair $(\mathcal{P},\mathcal{M})$ where $\mathcal{P}$ is a nonempty set of positions and $\mathcal{M}(p)\subseteq\mathcal{P}$ is the (finite) set of legal moves from $p$, with the property that the move relation does not depend on which player is to play (symmetry). We assume normal play: the player who makes the last move wins.
A position with no legal moves is terminal.
4.2 NIM
Definition 4.2 (NIM).
A NIM position consists of $h$ heaps with sizes $x_{1},\dots,x_{h}\in\mathbb{N}_{0}$. A legal move reduces exactly one heap: choose $i$ with $x_{i}>0$ and replace $x_{i}$ by some $x_{i}^{\prime}<x_{i}$.
4.3 Sprague–Grundy theory and nim-sum
The Sprague–Grundy theorem states that every finite impartial game position is equivalent (under disjunctive sum) to a NIM heap of some size (its nimber/Grundy value). For NIM itself, the Grundy value of a position is the bitwise XOR (nim-sum) of its heap sizes.
Definition 4.3 (Nim-sum).
For heap sizes $x_{1},\dots,x_{h}$, define
$\operatorname{nim}(x_{1},\dots,x_{h})\;=\;x_{1}\oplus x_{2}\oplus\cdots\oplus x_{h},$
where $\oplus$ denotes bitwise XOR.
Theorem 4.4 (Classical NIM characterisation).
A NIM position $(x_{1},\dots,x_{h})$ is losing for the player to move iff $\operatorname{nim}(x_{1},\dots,x_{h})=0$. Moreover, from any winning position (nim-sum $\neq 0$) there is a move to a losing position (nim-sum =0=0).
Why NIM is the canonical barrier.
The Sprague–Grundy theorem implies that mastering any impartial game reduces to two tasks: (1) computing the Grundy values (nimbers) of sub-components, and (2) computing their nim-sum. Our impossibility results target the second step. By proving that fixed-scale networks cannot compute the nim-sum (Theorem 5.2), we establish a fundamental barrier for the entire class of impartial games: no matter how easily an agent might perceive the local structure of a game (e.g., in Kayles or Dawson’s Kayles), it cannot aggregate these local values into a global decision whenever the aggregation logic relies on parity. Thus, NIM is not merely a specific game, but the isolate computation required to solve combinatorial sums of games.
5 Main results: impossibility and possibility
We separate the negative results for single-frame representations from the positive results obtained by adding recent history.
5.1 Mastery notions
We distinguish three notions of mastery:
Definition 5.1 (Strong, weak, and restoration mastery).
An agent achieves strong mastery of NIM if, for every reachable position $p$ with the agent to move, it plays optimally (i.e., implements a winning strategy whenever one exists).
It achieves weak mastery relative to an initial position $p_{0}$ if it plays optimally on every position that is reachable from $p_{0}$ under some sequence of legal moves, against any opponent strategy.
It achieves restoration mastery if, whenever the opponent moves from a $P$-position $Q$ with $\operatorname{nim}(Q)=0$ to a position $Q^{\prime}$, the agent responds with a move $Q^{\prime}\to R$ such that $\operatorname{nim}(R)=0$.
We use standard impartial-game terminology: a $P$-position is losing for the player to move, and an $N$-position is winning. For NIM, Theorem 4.4 says that $P$-positions are exactly the positions with nim-sum 0.
5.2 Single-frame impossibility results
In a single-frame representation, the network sees only the current position. To play optimally, it must in effect determine whether the nim-sum is zero and, if not, find a move producing nim-sum zero. The core obstacle is that global nim-sum computation requires parity-like behaviour across many bits, and parity is not in $\mathrm{AC}^{0}$ [5, 7, 24, 18].
Theorem 5.2 (Single-frame strong mastery requires parity).
Consider the restricted family of NIM positions $P(\mathbf{b})=(2,b_{1},\dots,b_{m})$ where $b_{i}\in\{0,1\}$. Any single-frame agent that achieves strong mastery must compute the parity $b_{1}\oplus\cdots\oplus b_{m}$. Since $\mathrm{PARITY}\notin\mathrm{AC}^{0}$, no polynomial-size, constant-depth fixed-scale quantised network can strongly master NIM.
Proof.
Let $s=\operatorname{nim}(2,b_{1},\dots,b_{m})=2\oplus(b_{1}\oplus\cdots\oplus b_{m})$. The most significant bit of $s$ is $2$ (the $2^{1}$ bit), and among the heaps only the first heap has that bit set. Therefore, any move that makes the nim-sum zero must change the first heap. The unique choice that makes the new nim-sum zero is
$x_{1}^{\prime}\;=\;2\oplus s\;=\;b_{1}\oplus\cdots\oplus b_{m}\in\{0,1\},$
so the winning move is uniquely determined by the parity of $(b_{1},\dots,b_{m})$.
Moreover, the Boolean input length $n$ of this family satisfies $n=\Theta(m)$ under any fixed-width binary encoding of the heaps, so this yields a standard $\mathrm{AC}^{0}$ hardness reduction from PARITY. ∎
5.3 Multi-frame possibility results
The fundamental idea is to avoid computing the global nim-sum from scratch. Instead, we exploit that each move changes only one heap, so the nimber difference between consecutive positions is locally computable.
Definition 5.3 (Nimber difference).
Given consecutive positions $P$ and $P^{\prime}$ in a play sequence, define the nimber difference
$\Delta(P,P^{\prime})\;=\;\operatorname{nim}(P)\ \oplus\ \operatorname{nim}(P^{\prime}).$
In NIM, if $P^{\prime}$ is obtained from $P$ by changing heap $i$ from $x_{i}$ to $x_{i}^{\prime}$, then
$\Delta(P,P^{\prime})\;=\;x_{i}\oplus x_{i}^{\prime},$
a purely local computation on the changed heap.
Lemma 5.4 (Computing nimber difference in $\mathrm{AC}^{0}$ for bounded changes).
Fix a constant $t\in\mathbb{N}$. Let $P$ and $P^{\prime}$ be NIM positions whose heap vectors differ in at most $t$ heaps. Then $\Delta(P,P^{\prime})$ can be computed by an $\mathrm{AC}^{0}$ circuit of polynomial size (in the input length).
Proof.
See Appendix E. ∎
5.3.1 Two-frame local restoration rule
The next observation is that if, after our move, we make the opponent face a zero nim-sum position, then the opponent’s next move produces some nimber difference $\delta$. If we respond with a move that produces the same nimber difference $\delta$, we restore nim-sum zero.
Proposition 5.5 (Local restoration principle).
Suppose the opponent moves from a position $Q$ with $\operatorname{nim}(Q)=0$ to a position $Q^{\prime}$ (so it is now our turn). Let $\Delta(Q,Q^{\prime})=\operatorname{nim}(Q)\oplus\operatorname{nim}(Q^{\prime})$ denote the nimber difference caused by the opponent’s move. If we respond with a move $Q^{\prime}\to R$ such that
$\Delta(Q^{\prime},R)\;=\;\Delta(Q,Q^{\prime}),$
then $\operatorname{nim}(R)=0$.
Proof.
By definition, $\operatorname{nim}(Q^{\prime})=\operatorname{nim}(Q)\oplus\Delta(Q,Q^{\prime})$. Since $\operatorname{nim}(Q)=0$ (by hypothesis), we have $\operatorname{nim}(Q^{\prime})=\Delta(Q,Q^{\prime})$. The player’s move creates a new position $R$ with $\operatorname{nim}(R)=\operatorname{nim}(Q^{\prime})\oplus\Delta(Q^{\prime},R)$. Substituting the condition $\Delta(Q^{\prime},R)=\Delta(Q,Q^{\prime})$:
$\operatorname{nim}(R)\;=\;\operatorname{nim}(Q^{\prime})\oplus\operatorname{nim}(Q^{\prime})\;=\;0.$
Thus, matching the opponent’s difference restores the invariant. ∎
Moreover, a concrete restoration move can be selected in $\mathrm{AC}^{0}$ from $(Q,Q^{\prime})$ by computing $\delta=\Delta(Q,Q^{\prime})$ and then choosing (with a fixed tie-break rule) a heap $i$ such that $x_{i}^{\prime}=x_{i}\oplus\delta<x_{i}$ (Appendix C).
[…]
Figure 2: A worked restoration example (three heaps). Starting from a $P$-position $Q$ (green), the opponent changes heap 2 ($5\to 4$), creating a nimber difference $\delta=1$. The agent instantly restores the $P$-position by changing heap 1 ($3\to 2$), which imposes the matching difference $3\oplus 2=1$.
Interpretation (“mirror the delta”).
Once we have made the opponent face a $P$-position, we do not need to recompute $\operatorname{nim}(\cdot)$ from scratch. We only need to mirror the observed change $\Delta(Q,Q^{\prime})$ in our response move. Figure 2 illustrates this mechanism on a concrete three-heap instance.
Proposition 5.5 is structural: it describes a response pattern that keeps returning the game to nim-sum zero, provided the player can (i) compute $\delta=\Delta(Q,Q^{\prime})$ from the last two frames and (ii) choose a legal move $Q^{\prime}\to R$ that realises the same $\delta$. In NIM, when the opponent changes exactly one heap, $\delta$ is simply $x_{i}\oplus x_{i}^{\prime}$, and testing whether a candidate response move realises $\delta$ is a local check on the heap being changed.
The significance for learnability is representational. A two-frame policy network can condition its response on $(Q,Q^{\prime})$ and hence on the locally computable $\delta$, while a single-frame policy must reconstruct the relevant global information from $Q^{\prime}$ alone. In the experimental sections we use standard Monte Carlo tree search and ordinary simulated games; the multi-frame advantage comes from what the policy/value networks can infer from the input, rather than from imposing special constraints on simulation trajectories.
5.3.2 Implementation overview
An $\mathrm{AC}^{0}$ implementation of the two-frame delta and restoration components can be organised into:
  • a binary encoding scheme for heap vectors (Appendix C),
  • a constant-depth detection of which heaps changed between frames,
  • an $\mathrm{AC}^{0}$ circuit computing $\Delta$ on those changed heaps (Lemma 5.4),
  • an $\mathrm{AC}^{0}$ selection mechanism for a response move realising a target $\delta$.
5.4 Beyond evaluation: search, rollouts, and the structural gap
Theorem 5.2 shows that no $\mathrm{AC}^{0}$ circuit can evaluate the nim-sum in a single forward pass. A natural follow-up is whether planning—testing candidate moves and simulating play, as in Monte-Carlo Tree Search (MCTS)—can compensate.
We analyse a stylised rollout interface: a planning algorithm queries a rollout policy $\pi$ by simulating a complete game in which both players follow $\pi$ from a start position, and it observes only the terminal win/loss bit. This abstraction isolates the information that a weak policy can expose through repeated simulation; it is not a claim about the optimisation dynamics of AlphaZero-style training (cf. Section 1.2).
Definition 5.6 (Rollout policy and rollout outcome).
A rollout policy is a (deterministic) function $\pi\colon\{0,1\}^{NB}\to\{\text{legal moves}\}$ mapping a NIM position to a legal move. We say $\pi$ is $\mathrm{AC}^{0}$-bounded if it is computed by a circuit of constant depth and size $\mathrm{poly}(NB)$.
A rollout from position $P$ under policy $\pi$ is the terminal play obtained by having both players follow $\pi$ from $P$. The rollout outcome is $\mathrm{RO}_{\pi}(P)\in\{0,1\}$, where $1$ means the first player wins.
Definition 5.7 (Single-policy and multi-policy-head agents).
A single-policy agent $\mathcal{A}=(E,\pi,S)$ consists of an $\mathrm{AC}^{0}$ value head $E$, a single $\mathrm{AC}^{0}$ rollout policy $\pi$, and a search budget $S$ (number of rollout queries). Each rollout query simulates a complete game in which both players follow $\pi$ from a specified start position, returning a single win/loss bit.
A $k$-policy-head agent $\mathcal{A}=(E,\pi_{1},\dots,\pi_{k},S)$ has $k$ distinct $\mathrm{AC}^{0}$ rollout policies. For each of its $S$ rollout queries, the agent selects a policy head $\pi_{j}$ and a start position, then observes the terminal outcome of a complete game played under $\pi_{j}$. The choice of which head to deploy may depend adaptively on the outcomes of previous queries.
Remark 5.8 (Why multiple heads are not redundant).
Since $\mathrm{AC}^{0}$ is closed under parallel composition, one could compute all $k$ policies simultaneously as a single $\mathrm{AC}^{0}$ circuit. The distinction is not about computational power at the circuit level but about the rollout interface: each rollout query simulates a complete game under one policy, and at each game step only that policy determines the next move. A single rollout trajectory under a fixed policy $\pi$ produces one outcome bit; merging $k$ circuits into a single circuit does not change the fact that any one rollout follows a single deterministic rule, yielding a single bit of information. Multiple heads are useful because the agent can issue separate rollout queries under different policies $\pi_{j}$, producing different game trajectories that reveal independent linear functionals of the nim-sum (cf. Propositions 5.9 and 5.10).
Finite-state shootout controller (FSM over policy heads).
Definition 5.7 allows the choice of rollout head to depend adaptively on the outcomes of previous rollout queries. To make this adaptivity explicit (without changing the $\mathrm{AC}^{0}$ constraints on each head), it is convenient to view head-selection as a small finite-state controller that sits outside the heads and only sees the terminal win/loss bit of each rollout.
Formally, fix rollout heads $\pi_{0},\dots,\pi_{k-1}$. A finite-state shootout controller is a tuple
$\mathcal{C}=(\mathcal{S},s_{\mathrm{init}},g,\delta,\mathsf{halt}),$
where $\mathcal{S}$ is a finite set of controller states, $s_{\mathrm{init}}\in\mathcal{S}$ is the initial state, $g\colon\mathcal{S}\to\{0,\dots,k-1\}$ selects which head to use for the next rollout query, $\delta\colon\mathcal{S}\times\{0,1\}\to\mathcal{S}$ updates the controller state after observing a rollout outcome bit $r=\mathrm{RO}_{\pi_{g(s)}}(P)$, and $\mathsf{halt}\subseteq\mathcal{S}$ is a set of halting states associated with an output decision. (When $\mathsf{halt}=\emptyset$, the controller simply runs for a fixed number of queries and outputs a decision at the end.)
This viewpoint does not increase the per-step computational power of the rollout heads: each $\pi_{j}$ remains $\mathrm{AC}^{0}$-bounded. Instead, the controller organises which $\mathrm{AC}^{0}$ head is iterated in a shootout and when to stop querying. Figure 3 depicts a particularly simple controller that turns the $B$ structured rollout heads from Proposition 5.9 into a sequential “bit-query” procedure for deciding whether $\sigma(P)=\operatorname{nim}(P)$ is zero (and hence whether $P$ is a $P$-position).
[…]
Figure 3: A minimal finite-state shootout controller for deciding whether $\sigma(P)=\operatorname{nim}(P)$ is zero using the $B$ rollout heads of Proposition 5.9. The controller queries heads sequentially and halts as soon as it can certify $\sigma(P)\neq 0$. Reaching the final transition to $\sigma(P)=0$ implicitly means $r_{0}=\cdots=r_{B-1}=0$.
Integration (channels vs. protocol). It is helpful to reinterpret Part (ii) of Theorem 5.14 in an “information-channel” language: each policy head $\pi_{j}$ defines a distinct rollout measurement channel which, when iterated to termination, returns a single bit $r_{j}=\mathrm{RO}_{\pi_{j}}(P)$ that depends on the hidden $B$-bit invariant $\sigma(P)$ (cf. the identities in Proposition 5.9). With a single head, repeated rollouts merely resample the same linear functional $\ell(\sigma)$ (Proposition 5.10), so additional search budget cannot reveal the missing bits. With multiple heads, the agent has access to multiple independent channels and can reconstruct $\sigma(P)$ from their outcomes (Proposition 5.9). The finite-state shootout controller introduced above is the natural generalisation of this idea: it does not change what any head can compute, but organises the heads into an adaptive querying protocol—the controller decides which channel to query next (and when to stop, or when to re-query for robustness) based on previously observed rollout outcomes. In this view, multi-head provides the channels, while the FSM provides the sequential control that turns those channels into a practical evaluation procedure.
Depth amplification.
Although a single $\mathrm{AC}^{0}$ circuit cannot compute parity, iterating an $\mathrm{AC}^{0}$ policy for $L$ game steps yields an effective depth $\Theta(L)$ computation, which can reveal parity information. The next proposition makes this explicit by exhibiting $B$ simple $\mathrm{AC}^{0}$ rollout heads whose outcomes recover all $B$ bits of the nim-sum.
Proposition 5.9 (Multi-policy-head rollout sufficiency).
A $B$-policy-head $\mathrm{AC}^{0}$ agent achieves strong mastery of NIM on $N$ heaps of bit-width $B$, using total computation $O(BN^{2}\cdot 4^{B})$. For $B=O(\log N)$ this is polynomial in $N$.
Proof outline.
For each $j\in\{0,\dots,B-1\}$ define a deterministic rollout policy $\pi_{j}$ that repeatedly subtracts $2^{j}$ from the leftmost heap of size at least $2^{j}$ (and otherwise subtracts $1$ from the leftmost nonzero heap). A rollout under $\pi_{j}$ has a well-defined total number of moves whose parity is
$\mathrm{RO}_{\pi_{0}}(P)=\sigma_{0},\qquad\mathrm{RO}_{\pi_{j}}(P)=\sigma_{j}\oplus\sigma_{0}\quad(j\geq 1),$
where $\sigma=\operatorname{nim}(P)$ and $\sigma_{j}$ denotes its $j$th bit. The search controller first queries $\pi_{0}$ to obtain $\sigma_{0}$, then queries each $\pi_{j}$ ($j\geq 1$) to obtain $\sigma_{j}\oplus\sigma_{0}$, and recovers $\sigma_{j}=(\sigma_{j}\oplus\sigma_{0})\oplus\sigma_{0}$; thus $B$ rollouts under $B$ distinct heads recover all $B$ bits of $\sigma$. Testing candidate moves (by running the same $B$ rollouts from each successor position) identifies a move to nim-sum 0. A complete proof is in Appendix F. ∎
A single policy head yields only one bit.
The preceding result relies on having multiple rollout channels. With one fixed rollout head, repeated simulation does not reveal additional independent information about the nim-sum.
Proposition 5.10 (Single-policy information bound).
Fix $j$ and the rollout policy $\pi_{j}$ from the proof outline of Proposition 5.9. For any position $P$, the rollout outcome $\mathrm{RO}_{\pi_{j}}(P)$ is a fixed linear functional $\ell_{j}(\sigma(P))$ of the nim-sum bits (indeed, $\ell_{0}(\sigma)=\sigma_{0}$ and $\ell_{j}(\sigma)=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$). Moreover, if $P_{i}$ is obtained from $P$ by modifying one heap (as when evaluating candidate moves), then
$\mathrm{RO}_{\pi_{j}}(P_{i})=\ell_{j}(\sigma(P))\oplus c_{i},$
where the correction bit $c_{i}$ is computable from the known heap values and the chosen modification. Hence all rollout queries under a single head reveal copies of the same unknown bit $\ell_{j}(\sigma(P))$.
Proof.
The linearity claim follows from the explicit formulas above. If $P_{i}$ differs from $P$ by changing one heap from $h$ to $h^{\prime}$, then $\sigma(P_{i})=\sigma(P)\oplus(h\oplus h^{\prime})$ and $\ell_{j}(\sigma(P_{i}))=\ell_{j}(\sigma(P))\oplus\ell_{j}(h\oplus h^{\prime})$, where $\ell_{j}(h\oplus h^{\prime})$ is known from the candidate move. ∎
Remark 5.11 (Exponential signal decay under noisy rollouts).
Proposition 5.10 analyses deterministic rollout heads. In practice, a learned policy is noisy: at each step, it selects the intended move with probability $1-\varepsilon$ and errs otherwise. The terminal outcome of a rollout of length $L$ then depends on the parity of the number of correct moves along the path.
If per-step errors are approximately independent, the standard identity for parity of biased Bernoulli sums gives
$\left|\Pr[\text{first player wins}]-\tfrac{1}{2}\right|\;=\;\tfrac{1}{2}(1-2\varepsilon)^{L},$
which decays exponentially in the rollout length $L$. Even strong per-move accuracy (e.g. $\varepsilon=0.01$) yields a near-fair-coin signal once $L\gtrsim 50$.
Majority voting over $M$ independent rollouts does not remove this bottleneck: distinguishing a bias of $(1-2\varepsilon)^{L}$ from $\tfrac{1}{2}$ requires
$M=\Omega\!\bigl((1-2\varepsilon)^{-2L}\bigr),$
which is exponential in $L$. This quantifies the gap in Figure 4(a): a single noisy rollout head produces not merely limited information but exponentially weak information about the parity-governed outcome, regardless of search budget.
The independence assumption is an idealisation; correlated errors can change constants, but the qualitative message of exponential decay with rollout length remains.
Two-frame rollouts can act as universal verifiers.
The two-frame restoration rule from Section 5.3.1 leads to a rollout policy whose single outcome bit already decides whether a position is a $P$-position.
Definition 5.12 (Universal verifier).
A rollout policy $\pi^{*}$ is a universal verifier if $\mathrm{RO}_{\pi^{*}}(P)=0$ iff $\operatorname{nim}(P)=0$ (equivalently, $P$ is a $P$-position), when both players follow $\pi^{*}$. In other words, the first player loses the rollout precisely when the start position is losing for the player to move.
Proposition 5.13 (Restoration is a universal verifier in the rollout interface).
The two-frame restoration strategy $\pi^{*}$ of Proposition 5.5 is a universal verifier for NIM and is computable in $\mathrm{AC}^{0}$.
Proof.
In the deterministic rollout interface of Definition 5.6, each query is launched from a specified start transition. From that point onward, whenever one player induces a transition with nimber difference $\delta$, the other player can compute $\delta$ from two frames and apply Proposition 5.5 to return to nim-sum 0. Thus the key invariant is that the second player to act after departure from a $P$-position restores nim-sum 0.
If the queried start position satisfies $\operatorname{nim}(P)=0$, the first mover must leave nim-sum nonzero, and the second mover repeatedly restores nim-sum 0; hence the first mover loses and $\mathrm{RO}_{\pi^{*}}(P)=0$. If $\operatorname{nim}(P)\neq 0$, the verifier bit flips correspondingly in this rollout interface, yielding $\mathrm{RO}_{\pi^{*}}(P)=1$. The concrete move-construction details used in Theorem 5.14(iii) are given in Appendix G.
The policy uses only the two-frame difference signal from Lemma 5.4 and fixed tie-breaking, so it is $\mathrm{AC}^{0}$-computable. ∎
Summary theorem.
Combining these ingredients yields a clean separation between single-policy single-frame planning and the two bypasses (multiple policy heads or two-frame inputs). Figure 4 provides a visual summary of this rollout-level separation.
[…]
Figure 4: The Rollout Gap. (a) A single-frame policy acts as a random walk on the parity hypercube; the signal mixes rapidly to $0.5$ (noise). (b) A two-frame agent acts as a Universal Verifier, executing a deterministic restoration path that preserves the truth value of the move perfectly.
Theorem 5.14 (The structural gap in search complexity).
Consider NIM on $N$ heaps of bit-width $B\geq 2$.
  1. (i)
    Single-frame, single-policy limitation. Let $\mathcal{A}=(E,\pi_{j},S)$ be a single-frame $\mathrm{AC}^{0}$ agent using one rollout head $\pi_{j}$ from above. Then all $S$ rollout queries reveal at most one unknown bit $\ell_{j}(\sigma(P))$ of the $B$-bit nim-sum. Consequently, on a random $N$-position the probability that $\mathcal{A}$ identifies a winning move is at most $2^{-(B-1)}+\mathrm{negl}(N)$.
  2. (ii)
    Single-frame, multi-policy-head sufficiency. A $B$-policy-head $\mathrm{AC}^{0}$ agent achieves strong mastery with total computation $O(BN^{2}\cdot 4^{B})$.
  3. (iii)
    Two-frame, single-policy sufficiency. A two-frame $\mathrm{AC}^{0}$ agent using the restoration policy $\pi^{*}$ achieves strong mastery with total computation $O(N^{2}\cdot 4^{B})$ by testing candidate first moves via the universal verifier (accepting those with $\mathrm{RO}_{\pi^{*}}=0$, indicating a $P$-position) and then maintaining $\operatorname{nim}=0$ by restoration.
Proof outline.
Part (i) combines Proposition 5.10 with standard $\mathrm{AC}^{0}$ lower bounds implying negligible correlation with parity-type predicates [7, 24, 18]: since all rollout outcomes under a single head are determined by one unknown parity bit (plus known data), and the agent’s decision function is $\mathrm{AC}^{0}$ in the remaining inputs, the agent cannot distinguish among the $2^{B-1}$ nim-sum values consistent with that bit. Parts (ii) and (iii) follow from Propositions 5.9 and 5.13 respectively. A complete proof is in Appendix G. ∎
6 Learning impartial games with neural networks
The results above are primarily representational: they identify what information a bounded-depth, fixed-scale quantised model can or cannot compute from its input, and what additional interfaces (history or multiple rollout heads) make optimal play possible. This section connects the theory to the training protocols used in Section 7 and highlights the main modelling choice ($\mathrm{AC}^{0}$ vs. $\mathrm{TC}^{0}$).
6.1 Practical takeaways
  • Single-frame is brittle to parity. In the worst case, fixed-scale quantised bounded-depth inference cannot reliably represent the nim-sum or a winning move selector (Section 5.2).
  • Two bypasses. A two-frame input exposes the transition signal $\Delta(P,P^{\prime})$ and supports restoration (Section 5.3.1), while multiple rollout heads can recover the missing nim-sum bits through depth amplification (Section 5.4).
  • Optimisation remains nontrivial. Even when a solution is expressible in $\mathrm{AC}^{0}$, discovering it from sparse win/loss rewards can be difficult; our experiments therefore isolate a dense restoration supervision signal and separately report search-based performance.
6.2 Teacher-guided data as a controlled source of transitions
Let $\pi_{\mathrm{opt}}$ denote an optimal NIM policy (move to nim-sum 0 whenever possible). For a noise parameter $\varepsilon\in[0,1]$, define a noisy teacher $\pi_{T}^{\varepsilon}$ that plays according to $\pi_{\mathrm{opt}}$ with probability $1-\varepsilon$ and otherwise chooses a uniformly random legal move. Occasional random moves broaden state coverage and, crucially for two-frame agents, generate diverse transition pairs $(Q,Q^{\prime})$ from which the local difference signal $\Delta(Q,Q^{\prime})$ can be inferred.
6.3 Approximate majority, parity, and the $\mathrm{AC}^{0}$ abstraction
One might object that modern networks can implement majority-like decisions, suggesting $\mathrm{TC}^{0}$ (threshold circuits) as the “right” complexity class. Indeed, if one allows scale-adaptive thresholds/biases or fan-in dependent re-scaling, constant-depth threshold models can represent parity-type computations.
Our focus is a different deployment regime: fixed-latency inference with fixed-scale, bounded-range quantisation, where each unit’s threshold is $O(1)$ in units of the smallest nonzero weight. In this regime, each gate has a constant-size witness and collapses to a constant-width DNF, yielding an $\mathrm{AC}^{0}$ abstraction (Theorem 2.4). Constant-depth $\mathrm{AC}^{0}$ circuits can implement approximate counting [1], yet they have exponentially small correlation with parity [5, 7, 24, 18]. This aligns with the empirical difficulty of learning parities [26, 27, 3] and supports using $\mathrm{AC}^{0}$ as a conservative abstraction for fixed-scale quantised inference. Section 6.4 gives a complementary Fourier/spectral interpretation of this local-versus-global parity separation.
6.4 A spectral viewpoint: low-degree Fourier mass vs. parity
A useful complementary lens is Fourier analysis on the Boolean hypercube. Let $f:\{0,1\}^{n}\to\mathbb{R}$ and define the Walsh–Hadamard characters
$\chi_{S}(x)\;=\;(-1)^{\sum_{i\in S}x_{i}},\qquad S\subseteq[n].$
The set $\{\chi_{S}\}_{S\subseteq[n]}$ is an orthonormal basis with respect to the uniform distribution on $\{0,1\}^{n}$. The Fourier coefficients are
$\hat{f}(S)\;=\;\mathbb{E}_{x\sim U(\{0,1\}^{n})}\bigl[f(x)\chi_{S}(x)\bigr],\qquad\text{and}\qquad f(x)=\sum_{S\subseteq[n]}\hat{f}(S)\chi_{S}(x).$
We refer to $|S|$ as the Fourier degree (or “frequency”) of the coefficient $\hat{f}(S)$.
Definition 6.1 (Fourier tail / spectral bandwidth).
For $k\in\mathbb{N}$, define the degree->k>k Fourier tail
$\mathrm{Tail}_{>k}(f)\;=\;\sum_{S\subseteq[n]:\,|S|>k}\hat{f}(S)^{2}.$
We say $f$ is $(k,\varepsilon)$-spectrally band-limited if $\mathrm{Tail}_{>k}(f)\leq\varepsilon$.
Parity is a pure high-frequency character.
For any nonempty $S\subseteq[n]$, the parity function $\mathrm{PARITY}_{S}(x)=\chi_{S}(x)$ has Fourier expansion supported on a single set: $\widehat{\mathrm{PARITY}_{S}}(S)=1$ and $\widehat{\mathrm{PARITY}_{S}}(T)=0$ for $T\neq S$. Thus its spectral “frequency” is exactly $|S|$.
$\mathrm{AC}^{0}$ has little high-frequency mass.
A seminal theorem of Linial–Mansour–Nisan shows that $\mathrm{AC}^{0}$ functions have almost all Fourier mass on low degrees: for $f$ computed by an $\mathrm{AC}^{0}_{d}$ circuit of size $M$, the tail above degree $k$ decays rapidly with $k$ (see [8], and sharper bounds in [25]). In particular, for $M=\mathrm{poly}(n)$ and constant depth $d$, any $\mathrm{AC}^{0}$ function has exponentially small correlation with parity on $\omega(\log^{c}n)$ bits (for fixed $c$).
Local vs. global parity in NIM.
In our encoding, the $j$th bit of the nim-sum is the parity of the $j$th bits across all $N$ heaps, hence a degree-$N$ character. By contrast, the $j$th bit of the two-frame difference $\Delta(P,P^{\prime})$ depends only on the old and new $j$th heap-bit (NIM moves change one heap), hence is a degree-$2$ character. This spectral separation mirrors the single-frame barrier and the two-frame restoration bypass.
Connection to optimisation (informal).
Fourier degree also aligns with a common empirical observation in deep learning: gradient-based training tends to fit low-frequency components before high-frequency ones (“spectral bias” / “frequency principle”) [17, 28]. Since parity is a maximally high-frequency character on the hypercube, this provides a complementary optimisation-based explanation for why parity-type invariants are difficult to learn in practice even when they are representable.
6.5 From global invariants to local checks
The core representational point is that the global invariant $\operatorname{nim}(P)$ is hard to compute from a single frame, whereas the transition quantity $\Delta(P,P^{\prime})$ is locally computable from two frames (Lemma 5.4). Section 7.2 isolates this effect by training one-frame and two-frame models on the same restoration targets.
7 Experiments
We evaluate the multi-frame restoration hypothesis with a supervised restoration dataset from teacher–random transitions (Section 7.2) and a complementary single-frame multi-head shootout evaluation (Section 7.3). Unless otherwise stated, both one-frame (1F) and two-frame (2F) agents share the same action parameterisation and network architecture (Section 7.1).
7.1 Neural architecture and training objective
7.1.1 State and action representation
We use a fixed-width binary encoding of heap sizes with $k$ bits per heap. The single-frame input is the encoding of the current heap vector. The two-frame input concatenates the encodings of the previous and current heap vectors. When a previous frame is unavailable (e.g. on the first move of a game), we set the previous frame equal to the current position so that the input dimension is constant and the 2F policy remains well-defined.
The action space is parameterised as “set heap $i$ to value vv” for $i\in\{1,\dots,N\}$ and $v\in\{0,1,\dots,2^{k}-1\}$, giving $N\cdot 2^{k}$ action logits. Illegal actions (those with $v\geq x_{i}$) are masked before sampling or taking an argmax.
7.1.2 Network architecture
Both agents use the same MLP template with two heads: a policy head producing $N\cdot 2^{k}$ logits and a value head producing a scalar in [1,1][-1,1]. The two-frame network input has dimension $2Nk$; the single-frame network input has dimension $Nk$.
Component Type Dimensions Activation/Normalisation
Input State representation Two-frame: $2Nk$ bits (1F: $Nk$) None
Hidden 1 Fully Connected 1F: $Nk\to H$; 2F: $2Nk\to H$ LayerNorm + ReLU
Hidden 2 Fully Connected $H\to H$ ReLU
Policy Head Linear Output $H\to N\cdot 2^{k}$ Log-Softmax
Value Head Linear Output $H\to 1$ Tanh
Table 1: Neural network architecture used in the experiments. $H$ denotes the hidden width (a fixed hyperparameter).
7.1.3 Training targets and loss
For each training input (single-frame state or two-frame state-pair), the target is a soft distribution $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ over optimal restoration moves. The policy head is trained with cross-entropy to this target. The value head is retained for architectural consistency across agents, but the main reported restoration metric is policy correctness.
7.1.4 Evaluation criteria in this version
In this version we report two task-level notions of performance rather than full game-level Elo tournaments. Move-level restoration performance is measured by functional correctness: on a test position with nonzero nim-sum, a move is counted as correct if it produces a successor state with nim-sum zero. Positions with zero nim-sum have no winning move (Theorem 4.4); when reporting functional correctness we therefore condition on nonzero nim-sum positions. Position-level evaluation performance is measured by $P/N$ classification accuracy and $N$-position recall in FSM-controlled shootouts (Section 7.3).
7.2 Supervised restoration learning from teacher–random transitions
This experiment isolates the restoration response as a supervised learning problem. In normal-play impartial games, a restoration response is the move that returns play to a losing ($P$-)position after the opponent departs from it. In Nim, this is equivalent to selecting an optimal (winning) move from the post-opponent position: if the previous position had nim-sum 0, then after any opponent move the position is winning, and any optimal reply restores nim-sum 0.
The goal is to test the core theoretical prediction of Sections 5.25.3 in a controlled, low-noise setting: does access to the transition $(p,p^{\prime})$ (two frames) make restoration learnable, even when one-frame models fail? This benchmark builds on the observation that parity-type invariants pose a challenge for reinforcement learning in impartial games [31, 32], while explicitly providing the transition information needed by Proposition 5.5.
7.2.1 Game and representations
We consider Nim with $N=20$ heaps. Each heap size is represented with $k=4$ bits (little-endian), so heap values lie in $\{0,\ldots,15\}$. Moves are represented in a fixed action space of size $N\cdot 2^{k}=20\cdot 16=320$: an action chooses a heap index $i$ and a new heap value $v$ (masked to legal moves $0\leq v<h_{i}$).
A 1-frame (1F) policy receives only the current position $p^{\prime}$ (80-bit input), while a 2-frame (2F) policy receives the concatenation $(p,p^{\prime})$ of the previous and current positions (160-bit input). Both agents use the shared MLP architecture described in Section 7.1 (two hidden layers of width $H=256$).
7.2.2 Dataset generation (teacher–random–teacher)
We generate a dataset of size $D=10^{6}$ examples of the form
$(p,p^{\prime})\mapsto\pi_{\mathrm{opt}}(\cdot\mid p^{\prime}),$
where $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ is a soft target distribution over optimal moves from $p^{\prime}$.
Each example is generated as follows:
  1. 1.
    Sample a random winning Nim position $s$ (nim-sum $\neq 0$).
  2. 2.
    Apply an optimal teacher move from $s$ to obtain $p$. For optimal play in Nim, $p$ has nim-sum 0 (a $P$-position for the next player).
  3. 3.
    Apply a uniformly random legal move from $p$ to obtain $p^{\prime}$.
  4. 4.
    Define the label $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ as the uniform distribution over all optimal (winning) moves from $p^{\prime}$.
This construction ensures that each training input $(p,p^{\prime})$ corresponds to the canonical restoration setting: $p$ is a $P$-position, the random move produces a winning position $p^{\prime}$, and the label is the set of restoration moves back to a $P$-position.
Why teacher–random–teacher?
We do not sample arbitrary triples $(P,P^{\prime},a)$ uniformly at random, because the restoration claim is about a specific interaction regime: the opponent departs a $P$-position and the agent must restore it. The teacher step ensures $P$ is a genuine $P$-position reachable under optimal play, the random move injects diversity in the departure $P\to P^{\prime}$, and the final teacher move provides a correct restoration label for $P^{\prime}$.
Compared to earlier empirical studies of impartial games [31, 32], we “help” the learner in two ways: (i) we represent heap sizes in binary rather than as raw integers, and (ii) we train on perfectly labelled restoration targets (the full set of optimal moves), rather than requiring the agent to infer these targets from sparse win/loss outcomes.
We split the dataset into 80% train / 10% validation / 10% test.
7.2.3 Training and checkpoints
We train each model with supervised learning using cross-entropy to the soft target distribution. We use Adam with learning rate $10^{-3}$ and batch size 512 for 200 epochs. To study performance as a function of training, we save $C=10$ checkpoints per model at epochs $\{1,23,45,67,89,112,134,156,178,200\}$.
7.2.4 Evaluation metrics
We evaluate restoration performance on held-out data:
Restoration accuracy (validation/test).
Given a held-out example, we decode the model’s greedy action (argmax over legal moves). We count a prediction as correct if the chosen action lies in the support of $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ (i.e., it is one of the optimal restoration moves). This accounts for the fact that multiple optimal moves may exist.
7.2.5 Results
One-frame does not learn restoration; two-frame eventually does.
Figure 5 shows validation restoration accuracy as a function of epoch. The 1F model remains near chance throughout training, stabilising around 0.079 (min 0.075, max 0.081 across checkpoints on the test set). Here “chance” refers to a random-legal-move baseline on the same test distribution (single-digit restoration accuracy, since only a small fraction of legal moves are restoring moves). In contrast, the 2F model learns the restoration mapping and reaches essentially perfect accuracy with enough training. On the test set (Figure 6), the 2F checkpoints achieve 0.209 at epoch 1, 0.538 at epoch 23, 0.876 at epoch 45, 0.989 at epoch 67, and 0.99998 by epoch 178 (remaining $\geq 0.99993$ at epoch 200).
Figure 5: Validation restoration accuracy vs. epoch on the restoration dataset ($N=20$, $k=4$, $D=10^{6}$). The 1F model stays near chance, while the 2F model exhibits a characteristic two-stage learning curve with an intermediate plateau around 0.54 before reaching near-perfect accuracy.
Figure 6: Test restoration accuracy of checkpointed agents ($N=20$, $k=4$, $D=10^{6}$). Checkpoints are saved at epochs $\{1,23,45,67,89,112,134,156,178,200\}$.
A learnability “plateau” explained by approximate-majority shortcuts.
The 2F learning curve is not monotone: it first rises to an intermediate plateau near $0.54$ and remains there for many epochs before “breaking through” to perfect restoration. We emphasise that this is a hypothesis consistent with our data distribution; see Section 6.3 for the circuit-complexity perspective on approximate majority. This plateau is consistent with an approximate-majority shortcut on the least significant parity bit: in our generated data, the least significant bit of the post-random nim-sum is slightly imbalanced (empirically $P[\mathrm{LSB}(\operatorname{nim}(p^{\prime}))=1]\approx 0.5355$), so a policy that predicts this bit as a constant can achieve accuracy around $0.54$ while still learning the easier aspects of the restoration response. Only later does training force the model to represent the full parity structure needed for exact restoration, at which point accuracy quickly approaches 1. This behaviour reflects the broader theme of Section 6.3: approximate majority-type correlations are $\mathrm{AC}^{0}$-compatible and easy to pick up, while exact parity requires sharper global consistency.
Overall, these results support the theoretical narrative: single-frame models fail to learn the parity-governed optimal response even under perfect supervision, while two-frame models can learn the restoration rule from transition data. Moreover, once a reliable restoration prior is learned, the transition-based representation yields strong and stable restoration performance. The configuration reported here ($N=20$, $k=4$, $D=10^{6}$) is representative of a broader experimental campaign. We also ran the 1F/2F comparison across different numbers of heaps, bit-widths, dataset sizes, and hyperparameter settings (hidden width, learning rate, training epochs). The qualitative pattern is consistent throughout: single-frame models remain near chance while two-frame models learn restoration, with the plateau and breakthrough dynamics described above. Code and data for all configurations are available at https://github.com/SR123/NIM.
7.3 Single-frame evaluation via FSM-controlled shootouts (multi-head)
Purpose.
We complement the two-frame restoration study (Section 7.2) with an experiment targeting the single-frame bypass from Proposition 5.9: multiple $\mathrm{AC}^{0}$-bounded rollout heads combined with a finite-state shootout controller (Figure 3). The goal is to test the rollout interface itself (how much information can be extracted by controlled shootouts), not to claim that self-play discovers the structured heads without additional inductive bias.
Setup.
We consider NIM with $N=20$ heaps and bit-width $k=4$ (so $B=k=4$ heads), matching the encoding of Section 7.1.1. We evaluate on a balanced test set of 2000 positions with $50\%$ $P$-positions ($\sigma=0$) and $50\%$ NN-positions ($\sigma\neq 0$).
Training the rollout heads.
We train a shared-trunk network with $B=4$ policy heads by imitation learning against the structured policies $\pi_{0},\dots,\pi_{B-1}$ from Proposition 5.9. Each head $\pi_{j}$ subtracts $2^{j}$ from the leftmost heap of size at least $2^{j}$ (falling back to subtracting 1 from the leftmost nonzero heap). When multiple heaps qualify for the same phase of $\pi_{j}$ (e.g. several heaps have $h_{i}\geq 2^{j}$), any choice among them produces a rollout with the same total move count and hence the same parity outcome. We therefore train with soft targets: for each position and head $j$, the label is uniform over all equivalent actions, and the loss is cross-entropy against this soft distribution. This avoids penalising the model for an arbitrary tie-breaking choice among equally valid heaps.
The network uses the same shared-trunk architecture as Table 1 (two hidden layers of width $H=256$ with LayerNorm), with $B$ independent linear output heads each producing $N\cdot 2^{k}=320$ action logits. Training uses Adam (learning rate $10^{-3}$, batch size 1024) for 12 epochs on $2\times 10^{5}$ imitation examples.
Rollout evaluation.
Each rollout is deterministic: both players follow the same greedy policy head $\pi_{j}$ (argmax over legal actions) until termination. The rollout outcome is 1 if the first player wins (makes the last move) and 0 otherwise. Since rollouts are deterministic, a single rollout per (position, head) pair suffices.
The per-head accuracy metric is whether the model’s rollout outcome matches the teacher’s rollout outcome—the correct channel bit—not whether it matches the $P/N$ label directly. Each head $j$ extracts a specific linear functional of the nim-sum ($r_{0}=\sigma_{0}$; $r_{j}=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$), and only the FSM combining all $B$ heads reconstructs the full $P/N$ classification.
FSM-controlled shootout evaluation.
Given a test position $P$, the evaluator runs the controller of Figure 3: it queries head $\pi_{0}$ first; if $r_{0}=1$ it immediately declares $\sigma(P)\neq 0$ (N-position). Otherwise it queries heads $\pi_{1},\pi_{2},\ldots,\pi_{B-1}$ sequentially, declaring N-position as soon as any $r_{j}=1$, or declaring P-position if all $B$ bits are 0. Early stopping reduces the average number of rollouts below $B$ on N-positions. We report (i) overall $P/N$ evaluation accuracy and (ii) mean number of shootouts used.
Results.
Table 2 reports the results.
Multi-head FSM achieves perfect classification. With all $B=4$ heads, the FSM controller correctly classifies every test position (1.000 accuracy), confirming Proposition 5.9: the $B$ structured rollout heads recover all $B$ bits of the nim-sum, and the FSM reconstructs the full $P/N$ classification. The learned model matches the teacher exactly, indicating that the soft-target imitation training produces heads whose rollout trajectories preserve the parity structure needed to extract the correct channel bits.
Single-head baseline and theoretical prediction. The single-head baseline achieves 0.768 overall accuracy and 0.536 N-position recall. These numbers have an exact theoretical explanation. With $k=4$ bits, an N-position has nim-sum $\sigma\in\{1,\dots,15\}$. Of these 15 values, exactly 8 are odd ($\sigma_{0}=1$: the values $1,3,5,7,9,11,13,15$) and 7 are even ($\sigma_{0}=0$). Head $\pi_{0}$ detects an N-position only when $\sigma_{0}=1$, so the theoretical N-position recall is $8/15\approx 0.533$ (observed: 0.536 on the finite test set). All P-positions are correctly classified (since $\sigma=0$ implies $\sigma_{0}=0$), so on a balanced test set the overall accuracy is $\tfrac{1}{2}\cdot 1+\tfrac{1}{2}\cdot\tfrac{8}{15}=\tfrac{23}{30}\approx 0.767$ (observed: 0.768). This is a quantitative instance of the single-policy bottleneck (Theorem 5.14(i)): one rollout head reveals only $\sigma_{0}$, leaving B1=3B-1=3 bits unresolved.
Connection to the two-frame restoration plateau. The $8/15$ ceiling is closely related to the $\approx 0.54$ accuracy plateau in the two-frame restoration experiment (Section 7.2, Figure 5). Both reflect the information limit of $\sigma_{0}$ alone: in the restoration setting, a model that has learned to exploit the least significant bit—but not the higher bits—can correctly restore only the fraction of positions determined by that single bit, plateauing near $0.54$. The FSM single-head recall of $8/15\approx 0.533$ is the same limitation viewed from the classification side. In both cases, the gap to perfect performance measures the missing B1=3B-1=3 higher nim-sum bits—information that the multi-head FSM recovers and the two-frame model eventually learns.
Early stopping. The average number of shootouts is 2.87 rather than the worst-case 4, because any N-position with $\sigma_{0}=1$ is caught at head 0 (one rollout); only those with $\sigma_{0}=0$ require further heads. Representativeness. The configuration reported here ($N=20$, $k=4$, $B=4$) is likewise representative. We ran the multi-head FSM evaluation for larger values of $N$ and $k$ (and correspondingly more heads) and varied hyperparameters including hidden width, learning rate, training set size, and number of epochs. The qualitative pattern is consistent: the multi-head FSM achieves perfect or near-perfect win/loss position classification whenever per-head imitation accuracy is sufficiently high, while the single-head baseline remains bounded by the $\sigma_{0}$-only ceiling. Code and data for all configurations are available at the same repository.
Method Heads Shootouts (avg.) $P/N$ Accuracy (Overall) NN-pos Recall
Single-head baseline (model) 1 1.00 0.768 0.536
Multi-head + FSM (model) 4 2.87 1.000 1.000
Single-head baseline (teacher) 1 1.00 0.768 0.536
Multi-head + FSM (teacher) 4 2.87 1.000 1.000
Table 2: Single-frame $P/N$ evaluation by FSM-controlled shootouts ($N=20$, $k=4$), reporting overall $P/N$ accuracy and $N$-position recall.
8 Discussion and broader implications
Our main message is that multi-frame representations can turn globally parity-governed decision rules into locally checkable transition rules, making them feasible for fixed-latency, fixed-scale quantised inference. We also show that a multi-policy-head architecture can break the parity barrier even in a single frame. We briefly discuss several directions where this perspective may be useful beyond the specific NIM setting.
8.1 Theoretical idealisation vs. practical architecture
The formal $\mathrm{AC}^{0}$ results describe a fixed-latency logical core with fixed-scale quantised thresholds. Our empirical networks additionally use LayerNorm and Softmax (implemented as Log-Softmax), both of which are statistical normalisation layers with global aggregation and therefore exceed strict $\mathrm{AC}^{0}$ as literal circuit operations. We treat these components as numerical stabilisation/calibration layers, distinct from the logical reasoning layers required to resolve parity. The key empirical pattern is consistent with this interpretation: in Section 7.2, the one-frame model still fails near chance despite using the same normalisation layers, while two-frame access succeeds once local transition information is exposed.
8.2 Multi-frame approaches to mathematical discovery
The $\mathrm{AC}^{0}$ perspective is not limited to game playing. It provides a general method for diagnosing when a direct neural approach is blocked by high-frequency signals (Section 6.4) and for designing alternative representations that bypass the barrier.
This principle guided our work on discovering large Condorcet domains [9, 30, 29]. The long-standing record domain (Fishburn) involves parity-like patterns and high-frequency signals across many triples; single-frame neural search struggled even at small parameters. Rather than using explicit multi-frame inputs, we used a hand-crafted AI-inspired search based on local database look-ups and locally checkable invariants. This transformed the search from an $\mathrm{AC}^{0}$-infeasible global constraint into tractable local constraints that guide exploration.
8.3 Beyond NIM: other impartial games
By Sprague–Grundy theory, every finite impartial game position has a nimber. However, the computational complexity of determining nimbers varies widely. Some impartial games are PSPACE-complete (e.g. Geography) [19]. In such cases, even polynomial-time algorithms (let alone $\mathrm{AC}^{0}$-bounded inference) cannot compute optimal play from arbitrary positions.
Our message is not that $\mathrm{AC}^{0}$ is universally sufficient, but that it is a useful model for diagnosing when learning failure is due to parity-like global constraints and for motivating representation/search strategies that reduce global constraints to local checks.
8.4 The multi-policy-head architecture and the single-policy bottleneck
Theorem 5.14 formalises an information bottleneck in our stylised deterministic rollout-policy interface (AlphaZero-like in spirit): with a single policy head, all rollouts under that policy reveal the same linear functional of the nim-sum (Proposition 5.10), so additional search does not recover the missing bits. A multi-policy-head architecture (Definition 5.7) provides multiple independent rollout channels; in NIM, $B$ suitably structured heads suffice to recover all $B$ bits of the nim-sum (Proposition 5.9).
This suggests a qualitative separation in search requirements: a single-policy agent must effectively brute-force missing information, whereas multiple heads can supply it directly. We leave formal training- and search-complexity bounds for future work.
More broadly, whenever optimal decisions depend on a multi-bit global invariant that is hard for $\mathrm{AC}^{0}$, adding independent policy heads is a principled architectural intervention suggested by circuit complexity.
8.5 The Optimisation Gap: Why Representation is Not Enough
Our theoretical results show that both the 2F architecture and the multi-policy-head architecture remove the representational barrier: solutions are implementable in $\mathrm{AC}^{0}$. The supervised results nevertheless indicate a complementary optimisation gap: even with the right representation, learning can pass through long intermediate plateaus before converging to exact restoration. In end-to-end self-play settings, where credit assignment is substantially noisier, this optimisation gap is expected to be even more pronounced; we therefore treat teacher-guided MCTS-style training as an important future direction rather than a completed empirical claim in this version.
While the 2F agent can represent the optimal restoration rule in $\mathrm{AC}^{0}$, discovering this rule via gradient descent from sparse rewards remains challenging. Similarly, while a multi-policy-head agent can represent the $B$ structured rollout policies, discovering them through self-play is an open problem. This distinction—between what an architecture can express and what training can find—defines the frontier of the research. We have established the necessary structural priors (history/difference access, or multiple policy heads), but future work must develop the algorithmic priors (e.g., difference-attention heads, curriculum design, auxiliary losses rewarding nim-sum-zero transitions) to make these solutions discoverable by standard RL optimisers.
8.6 Significance and impact
Our results suggest several broader implications.
  • For learning theory: sequential interaction can transform a globally hard invariant into a locally checkable transition rule, changing what weak learners can achieve. Multiple independent policy heads can extract multi-bit invariants that any single policy head provably cannot.
  • For neural architecture: short history windows (or difference features) can be a principled way to bypass parity-like barriers without scaling depth/width. When history is unavailable, multiple policy heads offer an alternative bypass within a single-frame architecture.
  • For practice: when a model fails, the cause may be structural misalignment rather than data scarcity; circuit complexity offers a diagnostic lens. The single-policy bottleneck identified here may affect AlphaZero-style systems in any domain governed by multi-bit global invariants.
9 Conclusion
We model fixed-latency neural inference in a fixed-scale, bounded-range quantised regime and show that such models are simulable by $\mathrm{AC}^{0}$ (Theorem 2.4). For NIM this yields a sharp single-frame barrier: optimal play depends on the global nim-sum, a parity-type invariant outside $\mathrm{AC}^{0}$.
We then show how interaction can bypass this barrier. Rollouts amplify effective depth, but with a single policy head every rollout reveals the same linear functional of the nim-sum (Proposition 5.10), so increased search budget does not recover the missing bits. Multiple policy heads break this bottleneck by providing independent rollout channels (Proposition 5.9). Alternatively, giving the agent two consecutive frames exposes the nimber difference $\Delta(P,P^{\prime})$, enabling a restoration policy that maintains $\operatorname{nim}=0$ and acts as a universal verifier (Propositions 5.5 and 5.13).
Our experiments support the restoration viewpoint in a controlled setting. On 20-heap, 4-bit NIM with $10^{6}$ supervised restoration examples, a one-frame policy stays near chance while a two-frame policy reaches near-perfect restoration accuracy (Section 7.2). In a complementary single-frame evaluation, a multi-head FSM shootout protocol reaches perfect win/loss position classification while the single-head baseline remains bottlenecked by one-bit information (Section 7.3). At the same time, the difficulty of end-to-end MCTS training highlights an optimisation gap: representational sufficiency does not by itself guarantee that standard RL optimisers will discover the solution from sparse rewards.
Overall, circuit complexity offers a concrete design principle for reliability beyond games. In domains such as logical verification, compiler optimisation, or symbolic reasoning, optimal decisions often hinge on global invariants (e.g. satisfiability or cycle structure) that can be hard for fixed-scale quantised bounded-depth inference. Our results suggest that rather than only scaling depth/width/search, a more effective bypass is to restructure inputs so local transition signals are explicit, converting a “state solving” problem into a “transition verification” problem.
Appendix A Formal definitions of fixed-scale quantised neural models
This appendix provides formal definitions of the fixed-scale quantised neural models used throughout the paper.
A.1 Preliminaries: fixed-scale quantisation and representable parameters
Fix constants $W,D\in\mathbb{N}$. Define the fixed finite set of representable parameters
$\mathcal{Q}_{W,D}=\Big\{\frac{\ell}{D}:\ell\in\{0,1,\dots,W\}\Big\}.$
In particular, the set of allowable weights/thresholds does not depend on input length $n$. This models fixed-scale quantisation with bounded range: values lie on a fixed grid (step $1/D$) and saturate at magnitude $W/D$, independent of the input length $n$. This is closer to deployment-time fixed-point/INT8-style inference with fixed scaling factors than to floating-point arithmetic. In particular, mere finiteness of a floating-point format is not what drives our $\mathrm{AC}^{0}$ containment; the key is that the effective range/scale of thresholds does not grow with fan-in as $n$ increases.
We will allow negations on wires (equivalently, $\mathrm{AC}^{0}$ has NOT gates). This is standard in circuit complexity and avoids conflating two independent issues: (i) parity hardness vs (ii) monotonicity restrictions.
A.2 Feedforward model
Definition A.1 (Fixed-scale quantised constant-depth neural network).
A fixed-scale quantised $\mathsf{NN}$ of depth $L$ on $n$ Boolean inputs is specified by:
  • Layer widths $m_{0}=n$, $m_{1},\dots,m_{L}$, where $m_{\ell}\leq\mathrm{poly}(n)$.
  • For each layer $\ell\in\{1,\dots,L\}$ and unit $i\in\{1,\dots,m_{\ell}\}$:
    • weights $w_{\ell,i,j}\in\mathcal{Q}_{W,D}$ for $j\in\{1,\dots,m_{\ell-1}\}$,
    • a threshold $\theta_{\ell,i}\in\mathcal{Q}_{W,D}$,
    • an optional negation bit $\eta_{\ell,i,j}\in\{0,1\}$ indicating whether input $j$ is negated before weighting.
The computation proceeds as follows. Let $y^{(0)}=x\in\{0,1\}^{n}$. For $\ell\geq 1$,
$y^{(\ell)}_{i}\;=\;\mathbf{H}\!\Big(\sum_{j=1}^{m_{\ell-1}}w_{\ell,i,j}\cdot(y^{(\ell-1)}_{j}\oplus\eta_{\ell,i,j})\;-\;\theta_{\ell,i}\Big),$
where $\mathbf{H}(z)=1$ if $z\geq 0$ and $\mathbf{H}(z)=0$ otherwise. The network output is $y^{(L)}$ (or a designated subset of its coordinates).
A.3 Recurrent model (finite window)
Definition A.2 (Fixed-scale quantised recurrent neural network with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-scale quantised $\mathsf{RNN}$ consists of:
  • input bits $x_{t}\in\{0,1\}^{n}$ at time $t$,
  • a hidden state $h_{t}\in\{0,1\}^{m}$ for some $m\leq\mathrm{poly}(n)$,
  • an output $o_{t}\in\{0,1\}^{k}$ for $k\leq\mathrm{poly}(n)$,
  • fixed-scale quantised parameters (weights/thresholds in $\mathcal{Q}_{W,D}$ and optional input negations) shared across time,
such that each update $h_{t+1}$ and $o_{t}$ is computed by a constant-depth fixed-scale quantised threshold network over inputs drawn from $\{x_{t^{\prime}},h_{t^{\prime}}:t-T<t^{\prime}\leq t\}$. Equivalently, unrolling for $T$ steps yields a constant-depth feedforward computation with depth $O(L\cdot T)$.
A.4 Attention-style model (finite window)
Definition A.3 (Fixed-scale quantised LTST model with finite window).
Fix a time window $T\in\mathbb{N}$. A fixed-scale quantised $\mathsf{LTST}$ model processes a sequence $(x_{t-T+1},\dots,x_{t})$ by:
  • computing short-term and long-term attention scores using fixed-scale quantised parameters in $\mathcal{Q}_{W,D}$,
  • combining attended summaries through a constant number of layers of fixed-scale quantised threshold computations,
  • producing outputs $o_{t}\in\{0,1\}^{k}$.
We assume that all attention coefficients, thresholds, and intermediate gating parameters are drawn from $\mathcal{Q}_{W,D}$ and that the number of attention heads/layers is constant, while the total number of units is polynomial in $n$. Unrolling the fixed window $T$ yields a constant-depth feedforward threshold computation of depth $O(L\cdot T)$.
Remark A.4 (Why these formal definitions match $\mathrm{AC}^{0}$).
Theorem 2.4 applies because $L,T,W,D$ are constants and each unit’s parameters come from a finite set independent of $n$. Appendix B gives an explicit circuit construction.
Appendix B An $\mathrm{AC}^{0}$ construction for simulating fixed-scale quantised networks
This appendix gives a clean circuit construction proving Theorem 2.4.
B.1 One threshold unit is in $\mathrm{AC}^{0}$ under fixed-scale quantisation
We first show that a single fixed-scale quantised threshold unit (as in Definition A.1) is computed by a depth-2 $\mathrm{AC}^{0}$ circuit of polynomial size.
Lemma B.1 (Fixed-scale quantised threshold gate $\in\mathrm{AC}^{0}$).
Fix constants $W,D\in\mathbb{N}$. Let $z_{1},\dots,z_{m}\in\{0,1\}$ be Boolean inputs, and let $w_{1},\dots,w_{m},\theta\in\mathcal{Q}_{W,D}$. Consider the Boolean function
$f(z_{1},\dots,z_{m})\;=\;\mathbf{H}\!\Big(\sum_{j=1}^{m}w_{j}z_{j}-\theta\Big).$
Then $f$ is computable by a depth-2 $\mathrm{AC}^{0}$ circuit of size $O(m^{K})$, where $K:=W$ is a constant (depending only on $W,D$).
Proof.
Multiply the inequality by $D$ to clear denominators. Write $w_{j}=a_{j}/D$ and $\theta=b/D$ with $a_{j}\in\{0,1,\dots,W\}$ and $b\in\{0,1,\dots,W\}$. Then
$f(z)=1\quad\Longleftrightarrow\quad\sum_{j=1}^{m}a_{j}z_{j}\;\geq\;b.$
If $b=0$ the function is identically $1$. Assume $b\geq 1$.
Because each $a_{j}$ is a nonnegative integer and each active input contributes at least $1$ to the left-hand side whenever $a_{j}\geq 1$, any satisfying assignment contains a subset of at most $b$ active variables whose coefficients already sum to at least $b$. More formally: if $\sum_{j}a_{j}z_{j}\geq b$, let $S=\{j:z_{j}=1\text{ and }a_{j}\geq 1\}$. Choose any multiset-free subset $T\subseteq S$ by greedily adding indices until the partial sum reaches $\geq b$. Since each added index increases the sum by at least $1$, this process stops after at most $b$ steps. Because $b\leq W$, we get $|T|\leq W=:K$.
Therefore,
$f(z)=1\quad\Longleftrightarrow\quad\bigvee_{\begin{subarray}{c}T\subseteq[m]\\ |T|\leq K\\ \sum_{j\in T}a_{j}\geq b\end{subarray}}\ \bigwedge_{j\in T}z_{j}.$
This is a depth-2 DNF: an OR of ANDs. The number of terms is at most $\sum_{r=0}^{K}\binom{m}{r}=O(m^{K})$, polynomial since $K$ is constant. Unbounded fan-in AND/OR gates give an $\mathrm{AC}^{0}$ circuit of depth $2$ and size $O(m^{K})$. ∎
Remark B.2 (Negations).
If some inputs are negated before weighting, i.e. $(z_{j}\oplus\eta_{j})$, this simply replaces $z_{j}$ by either $z_{j}$ or $\neg z_{j}$ in each conjunction term. Since $\mathrm{AC}^{0}$ allows NOT gates, Lemma B.1 still applies.
B.2 From one unit to a full network
Lemma B.3 (One layer of fixed-scale quantised units is in $\mathrm{AC}^{0}$).
Fix $W,D$ and let a layer consist of $m^{\prime}$ threshold units whose inputs are outputs of a previous layer of width $m$, with $m,m^{\prime}\leq\mathrm{poly}(n)$. Then the layer can be simulated by an $\mathrm{AC}^{0}$ circuit of constant depth and polynomial size.
Proof.
Apply Lemma B.1 independently to each unit. Each unit yields a depth-2 circuit of size $O(m^{K})$. Combine these $m^{\prime}$ circuits in parallel; the total size is $m^{\prime}\cdot O(m^{K})=\mathrm{poly}(n)$ and depth remains $2$. ∎
Proof of Theorem 2.4.
For a feedforward $\mathsf{NN}$ of constant depth $L$, repeatedly replace each layer by the $\mathrm{AC}^{0}$ circuit from Lemma B.3, wiring outputs of layer $\ell-1$ circuits as inputs to layer $\ell$ circuits. Depth multiplies by a constant factor: overall depth is $O(L)$ (constant), and size remains polynomial as a composition of polynomial-size layers.
For $\mathsf{RNN}$ and $\mathsf{LTST}$ with time window $T$, unroll the computation for $T$ steps (constant) to obtain a feedforward computation of depth $O(LT)$ (still constant). Apply the same layer-by-layer replacement. ∎
Appendix C Implementation details (overview)
This appendix summarises the components needed to implement the two-frame delta computation and the local restoration response within $\mathrm{AC}^{0}$ constraints.
  • Encoding. Represent each heap size in binary using a fixed block of bits; represent a position as the concatenation of blocks.
  • Difference detection. Given two frames $P,P^{\prime}$, compute in parallel which heap blocks differ (bitwise XOR within each block, then OR across the block).
  • Compute $\Delta$. If positions differ in at most $t$ heaps, compute the XOR of the changed-heap values (Appendix E). Since $t$ is constant, this is constant-depth.
  • Move selection. Given $\delta$ and current position, in parallel for each heap compute the candidate new size $x_{i}^{\prime}=x_{i}\oplus\delta$ and test legality ($x_{i}^{\prime}<x_{i}$). Use a fixed tie-breaking rule to output one legal move (e.g. smallest index).
All subroutines are implementable in $\mathrm{AC}^{0}$ because comparisons and equality on $O(\log m)$-bit blocks have constant-depth implementations (e.g., depth-3 with unbounded fan-in).
Appendix D Weak mastery examples
Example D.1 (Pairing strategy).
Consider positions $(a,a,b,b,c,c,\dots)$ consisting of pairs of equal heaps. The second player can mirror the first player’s move on the matching heap, maintaining the paired structure until termination, hence winning. This strategy is implementable by simple local equality checks and does not require global nim-sum computation.
Appendix E Proof of Lemma 5.4
We provide an explicit $\mathrm{AC}^{0}$ construction for computing nimber differences between positions differing in at most $t$ heaps.
Proof of Lemma 5.4.
Let positions be encoded as $h$ blocks of length $B$ bits (heap sizes in binary), so total input length is $n=hB$.
Step 1: identify changed heaps. For each heap $i$ and bit position $k\in[B]$, compute $d_{i,k}=x_{i,k}\oplus x^{\prime}_{i,k}$. Then $c_{i}=\bigvee_{k\in[B]}d_{i,k}$ indicates whether heap $i$ changed. This is depth-2 (XOR can be written as depth-2 using AND/OR/NOT on two bits; the OR over $B$ bits is unbounded fan-in).
Step 2: assume at most $t$ changes. By promise, $\sum_{i}c_{i}\leq t$ with $t$ constant.
Step 3: compute per-changed-heap XOR values. For each heap $i$, define the block XOR value $y_{i}=x_{i}\oplus x^{\prime}_{i}$ (bitwise XOR across the $B$ bits). Each output bit is XOR of two bits, hence $\mathrm{AC}^{0}$.
Step 4: XOR the at-most-$t$ nonzero blocks. Since at most $t$ heaps changed, $\Delta=\bigoplus_{i:c_{i}=1}y_{i}$ is XOR of at most $t$ $B$-bit numbers. For each bit position $k$, the $k$-th bit of $\Delta$ is parity of at most $t$ bits $\{y_{i,k}:c_{i}=1\}$. Because $t$ is constant, this parity can be computed by a constant-size DNF (or by hardwiring the truth table) and thus is in $\mathrm{AC}^{0}$.
The overall circuit has constant depth and polynomial size in $n$. ∎
Appendix F Additional proofs for the rollout analysis
Proof of Proposition 5.9.
Fix a position $P$ with heap sizes $h_{1},\dots,h_{N}\in\{0,\dots,2^{B}-1\}$ and let $\sigma=\operatorname{nim}(P)=h_{1}\oplus\cdots\oplus h_{N}$. For each $j\in\{0,\dots,B-1\}$ define the deterministic rollout policy $\pi_{j}$:
“Subtract $2^{j}$ from the leftmost heap of size at least $2^{j}$; if none exists, subtract $1$ from the leftmost nonzero heap.”
Each $\pi_{j}$ is $\mathrm{AC}^{0}$-computable because it performs constant-depth comparisons to the threshold $2^{j}$ and selects the leftmost qualifying heap.
Consider a rollout under $\pi_{j}$ from $P$. While some heap is at least $2^{j}$, every move subtracts exactly $2^{j}$ from some heap, so the number of such moves is $L_{1}=\sum_{i=1}^{N}\lfloor h_{i}/2^{j}\rfloor$. After this phase all heaps satisfy $h_{i}<2^{j}$, and $\pi_{j}$ falls back to subtracting $1$ until termination, for a further $L_{2}=\sum_{i=1}^{N}(h_{i}\bmod 2^{j})$ moves. Thus the total number of moves is $L=L_{1}+L_{2}$, and (normal play) the first player wins iff $L$ is odd, i.e. iff $L\bmod 2=1$.
For each $i$, the parity of $\lfloor h_{i}/2^{j}\rfloor$ equals the $j$th bit of $h_{i}$, denoted $h_{i}^{(j)}$, so $L_{1}\bmod 2=\bigoplus_{i}h_{i}^{(j)}=\sigma_{j}$. Similarly, $(h_{i}\bmod 2^{j})\bmod 2=h_{i}^{(0)}$, so $L_{2}\bmod 2=\bigoplus_{i}h_{i}^{(0)}=\sigma_{0}$. Hence $\mathrm{RO}_{\pi_{0}}(P)=\sigma_{0}$ and $\mathrm{RO}_{\pi_{j}}(P)=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$. From these $B$ outcomes the agent recovers $\sigma$.
To select a winning move, the agent enumerates all legal moves $m$ from $P$ (at most $N(2^{B}-1)$ possibilities), runs the $B$ rollouts from the successor position $P_{m}$, reconstructs $\sigma(P_{m})$, and accepts the first $m$ with $\sigma(P_{m})=0$. Each rollout has length $O(N2^{B})$, so the total computation is $O\bigl(N2^{B}\cdot B\cdot N2^{B}\bigr)=O(BN^{2}\cdot 4^{B})$. ∎
Appendix G Proof of Theorem 5.14
We prove each part of the theorem separately.
Proof of Theorem 5.14.
Part (i): Single-frame, single-policy limitation.
Let $\mathcal{A}=(E,\pi_{j},S)$ be a single-frame $\mathrm{AC}^{0}$ agent with value head $E$, a single rollout policy $\pi_{j}$, and search budget $S$. We show that $\mathcal{A}$ identifies a winning move on a uniformly random $N$-position with probability at most $2^{-(B-1)}+\mathrm{negl}(N)$.
Step 1: All rollout outcomes reduce to one unknown bit. By Proposition 5.10, the rollout outcome from any position $Q$ under $\pi_{j}$ satisfies $\mathrm{RO}_{\pi_{j}}(Q)=\ell_{j}(\sigma(Q))$, where $\ell_{j}\colon\mathbb{F}_{2}^{B}\to\mathbb{F}_{2}$ is a fixed $\mathbb{F}_{2}$-linear functional of the nim-sum $\sigma(Q)=\operatorname{nim}(Q)$.
During search the agent may test candidate moves $m$ from the root position $P$ by running rollouts from the successor positions $P_{m}$. Since $\sigma(P_{m})=\sigma(P)\oplus(h_{i}\oplus h_{i}^{\prime})$ for a move changing heap $i$ from $h_{i}$ to $h_{i}^{\prime}$, linearity gives
$\mathrm{RO}_{\pi_{j}}(P_{m})\;=\;\ell_{j}\bigl(\sigma(P)\bigr)\;\oplus\;\underbrace{\ell_{j}(h_{i}\oplus h_{i}^{\prime})}_{\text{known from move}~m},$
so every rollout outcome is a known affine function of the single unknown bit $\beta:=\ell_{j}(\sigma(P))$. Regardless of the budget $S$ and the choice of candidate moves, the agent learns at most the value of $\beta$.
Step 2: The agent’s total information. After all rollout queries the agent possesses:
  1. (a)
    the position $P$ (all $NB$ input bits),
  2. (b)
    the output $E(P)$ of the $\mathrm{AC}^{0}$ value head,
  3. (c)
    the single bit $\beta$.
Its decision function is therefore $f\colon\{0,1\}^{NB}\times\{0,1\}\to\{\text{moves}\}$. For each fixed $\beta\in\{0,1\}$, the map $P\mapsto f(P,\beta)$ is computable by an $\mathrm{AC}^{0}$ circuit (it composes $E$, the rollout policy, and Boolean operations on the known correction bits).
Step 3: One linear functional is insufficient for move selection. A winning move from an $N$-position $P$ requires finding a heap $i$ and setting $x_{i}^{\prime}:=x_{i}\oplus\sigma(P)$ with $x_{i}^{\prime}<x_{i}$. This depends on the exact $B$-bit value of $\sigma(P)\in\{1,\dots,2^{B}-1\}$. The bit $\beta=\ell_{j}(\sigma(P))$ imposes one $\mathbb{F}_{2}$-linear constraint, partitioning the nonzero nim-sum values into two fibres: one of size $2^{B-1}$ and one of size $2^{B-1}-1$ (since $\sigma=0$ belongs to exactly one fibre but is excluded for $N$-positions). Within a fibre, distinct values $\sigma\neq\sigma^{\prime}$ generically require different winning moves, because $x_{i}\oplus\sigma\neq x_{i}\oplus\sigma^{\prime}$ whenever $\sigma\neq\sigma^{\prime}$.
Step 4: Probability bound via $\mathrm{AC}^{0}$ correlation limits. Draw $P$ uniformly at random from $N$-positions with $N$ heaps of bit-width $B$. For large $N$ and fixed $B$, the conditional distribution of $\sigma(P)$ given $\beta$ is exponentially close (in $N$) to uniform over the relevant fibre. (This follows from standard concentration: the nim-sum is the XOR of $N$ independent $B$-bit values, and conditioning on one linear functional leaves the remaining B1B-1 bits nearly uniform.)
For each fixed $\beta$, the agent commits to a move via $f(P,\beta)$. Since $P\mapsto f(P,\beta)$ is $\mathrm{AC}^{0}$, the Linial–Mansour–Nisan theorem [8] and Håstad’s switching lemma [7] imply exponentially small correlation with any parity-type predicate on the $NB$ input bits. In particular, $f(\cdot,\beta)$ cannot distinguish among the $2^{B-1}-1$ possible nonzero nim-sum values consistent with $\beta$, except with advantage $\mathrm{negl}(N)$.
The probability of selecting a winning move is therefore at most
$\frac{1}{2^{B-1}-1}+\mathrm{negl}(N)\;\leq\;2^{-(B-1)}+\mathrm{negl}(N).$
Part (ii): Single-frame, multi-policy-head sufficiency.
This is Proposition 5.9, proved in full in Appendix F. The $B$ rollout heads $\pi_{0},\dots,\pi_{B-1}$ yield outcomes
$\sigma_{0}=\mathrm{RO}_{\pi_{0}}(P),\qquad\sigma_{j}=\mathrm{RO}_{\pi_{j}}(P)\oplus\sigma_{0}\quad(j\geq 1),$
from which the search controller recovers all $B$ bits of $\sigma(P)$. The agent then enumerates candidate moves $m$, runs the $B$ rollouts from each successor $P_{m}$, reconstructs $\sigma(P_{m})$, and accepts the first $m$ with $\sigma(P_{m})=0$. Total computation: $O(BN^{2}\cdot 4^{B})$.
Part (iii): Two-frame, single-policy sufficiency.
Let $\pi^{*}$ denote the two-frame restoration policy of Proposition 5.5, augmented with a fixed fallback rule for the opening move (where no previous-frame transition is available). By Proposition 5.13, $\pi^{*}$ is a universal verifier: $\mathrm{RO}_{\pi^{*}}(Q)=0$ iff $\operatorname{nim}(Q)=0$.
The agent proceeds from an $N$-position $P$ as follows.
  1. (i)
    Find a $P$-position successor. Enumerate all legal moves $m$ from $P$ (at most $N(2^{B}-1)$ candidates). For each, run a single rollout under $\pi^{*}$ from the successor position $P_{m}$. Accept the first move $m$ for which $\mathrm{RO}_{\pi^{*}}(P_{m})=0$, indicating $\operatorname{nim}(P_{m})=0$. Since $P$ is an $N$-position, Theorem 4.4 guarantees at least one such move exists, so the search terminates successfully.
  2. (ii)
    Maintain $\operatorname{nim}=0$ by restoration. After the initial move the opponent faces a $P$-position $P_{m}$. Whatever the opponent plays, the agent observes the two-frame transition $(P_{m},\,P_{m}^{\prime})$ and applies the restoration rule of Proposition 5.5: it computes $\delta=\Delta(P_{m},P_{m}^{\prime})$ and responds with a move that realises the same nimber difference $\delta$, returning to $\operatorname{nim}=0$. No further rollouts are needed; all subsequent moves use only the $\mathrm{AC}^{0}$-computable two-frame restoration.
The computation cost is dominated by the initial search: at most $N(2^{B}-1)$ candidate moves, each tested with one rollout of length at most $O(N\cdot 2^{B})$ steps, giving total computation $O(N^{2}\cdot 4^{B})$. ∎